### v1

# Model Internal Sleuthing: Finding Lexical Identity and Inflectional Morphology in Modern Language Models

Abstract
Large transformer-based language models dominate modern NLP, yet our understanding of how they encode linguistic information is rooted in studies of early models like BERT and GPT-2. To better understand today’s language models, we investigate how both classical architectures (BERT, DeBERTa, GPT-2) and contemporary large language models (Pythia, OLMo-2, Gemma-2, Qwen2.5, Llama-3.1) represent lexical identity and inflectional morphology. We train linear and nonlinear classifiers on layer-wise activations to predict word lemmas and inflectional features. We discover that models concentrate lexical information linearly in early layers and increasingly nonlinearly in later layers, while keeping inflectional information uniformly accessible and linearly separable throughout the layers. Further analysis reveals that these models encode inflectional morphology through generalizable abstractions, but rely predominantly on memorization to encode lexical identity. Remarkably, these patterns emerge across all 16 models we test, despite differences in architecture, size, and training regime (including pretrained and instruction-tuned variants). This consistency suggests that, despite substantial advances in LLM technologies, transformer models organize linguistic information in similar ways, indicating that these properties could be fundamental for next token prediction and are learned early during pretraining. Our code is available at https://github.com/ml5885/model_internal_sleuthing
Model Internal Sleuthing: Finding Lexical Identity and Inflectional Morphology in Modern Language Models
[…]
1 Introduction
Large transformer-based language models (LMs) are widely used for tasks such as text generation, question answering, and code completion (Workshop, 2023; Groeneveld et al., 2024; Grattafiori et al., 2024; Hui et al., 2024). However, how these models internally represent linguistic information remains an active research area. Prior work suggests a hierarchical organization where different layers specialize in capturing distinct levels of linguistic structure, from surface features to syntax and semantics (Jawahar et al., 2019; Tenney et al., 2019; Rogers et al., 2020).
But these studies focus only on first-generation LMs such as BERT and GPT-2 (Devlin et al., 2019; Radford et al., 2019). Since then, language technology has transformed dramatically - today’s models are far larger, trained on much more data, and adapted through extensive post-training procedures (Brown et al., 2020; Groeneveld et al., 2024; Lambert et al., 2025). We ask: how do modern LMs encode linguistic structure, and how do these representations differ from those in earlier models?
Consider the words walk, walked, jump, and jumped. Do language models group words by shared meaning (walk, walked) or by shared grammar (walked, jumped)? More broadly, where and how do LMs encode a word’s lemma (its lexical identity) and its grammatical form (inflectional morphology)?
To answer this, we train classifiers to predict either a word’s lemma or its inflection (e.g., tense, number, degree) from hidden state activations, using annotated sentences from the Universal Dependencies GUM corpus (Zeldes, 2017). We evaluate classifiers on 16 language models spanning diverse architectures and sizes. We find that:
  1. 1.
    Lexical information is predominantly encoded in early layers, while inflectional morphology is distributed more uniformly across layers.
  2. 2.
    These layer-wise distributions hold across all models, despite large-scale changes in architecture, amount of pretraining, and post-training procedures.
  3. 3.
    The effective dimensionality of representations generally declines with layer depth, though in some models the lowest point is in the middle layers - where just one or two dimensions capture nearly all the variance.
2 Model Internal Sleuthing
Figure 1: Overview of our classifier methodology. We extract hidden state activations from each model layer for target words and train two classifiers: one predicting inflection and another predicting lemma. We evaluate three classifier architectures (linear regression, MLP, and random forest) for each task to deduce information accessibility and linear separability across the layers of the language model. Performance differences between linear and non-linear classifiers reveal the encoding structure of linguistic information.
To study how language models encode linguistic information, we train simple probabilistic classifiers on the models’ internal representations or activations. [^1] We use representations, activations, and neurons interchangeably. By observing whether classifiers can recover word-level properties like lemma or inflection from a given representation, we can infer which layers, if any, linguistic information is present, and how directly (e.g., linearly or not). High classifier performance at a given layer suggests that the relevant information is encoded and accessible at that layer. Additionally, comparing performance across layers tells us how lexical and inflectional information is distributed throughout the model.
Figure 1 illustrates our approach. For each layer of each model, we extract hidden state activations for a target word and use them as input features to train specialized classifiers that predict either inflection or lemma. In our experiments, we train three different classifiers: linear regression, two-layer multi-layer perceptron (MLP), and random forest.
2.1 Linear Regression Classifier
Consistent with best practices for probing Hewitt and Liang (2019); Liu et al. (2019), we use simple linear regression classifiers. These models are solved via closed-form ridge regression Hastie et al. (2009), using the equation:
$W=(X_{\text{train}}^{T}X_{\text{train}}+\lambda I)^{-1}X_{\text{train}}^{T}Y_{\text{train}}$ (1)
where $X_{\text{train}}\in\mathbb{R}^{m\times d}$ contains $d$-dimensional representations for each of the $m$ examples in the training set, $Y_{\text{train}}\in\mathbb{R}^{m\times c}$ is the one-hot encoded matrix of training labels for $c$ classes, $\lambda$ controls the strength of L2 regularization, and $W\in\mathbb{R}^{d\times c}$ is the learned weight matrix of the classifier. Test predictions are given by:
$\hat{Y}_{\text{test}}=X_{\text{test}}W$ (2)
2.2 Random Forest Classifier
Random forests provide a complementary approach to linear models by handling complex feature interactions and decision boundaries. Unlike linear models that assume information is linearly separable, random forests combine an often large number of decision trees via model averaging to robustly identify patterns (Breiman, 2001).
Since our classification tasks involve a large number of classes (thousands of unique lemmas and multiple inflection categories), we use a one-vs-all approach. For each class $j$, we train a binary random forest classifier that distinguishes class $j$ from all other classes. Each binary classifier outputs a probability score $s_{j}(x)\in[0,1]$ representing the confidence that input $x$ belongs to class $j$. The final prediction is assigned based on the maximum score across all classes:
$\hat{y}=\operatorname{argmax}_{j\in\{1,2,...,c\}}s_{j}(x)$ (3)
where $s_{j}(x)$ is the probability that the trained binary classifier assigns to class $j$ for the input $x$ [^2] We use scikit-learn’s implementation (Pedregosa et al., 2011) with the predict_proba() method to obtain probability scores.
2.3 MLP Classifier
To test for more general non-linearity, we train a simple two-layer MLP with ReLU activation, defined as:
$\hat{Y}=\text{softmax}(\text{ReLU}(X_{\text{train}}W_{1})W_{2})$ (4)
where $W_{1}\in\mathbb{R}^{d\times h}$ and $W_{2}\in\mathbb{R}^{h\times c}$ are the weight matrices and $h=64$ is the hidden layer size (Rosenblatt, 1958) [^3] Bias terms are omitted for brevity. Two-layer MLPs with ReLU activation are universal function approximators, capable of approximating any continuous function to arbitrary precision given sufficient width (Hornik et al., 1989).
Comparing performance across these three classifier types provides insights into how information is encoded: if linear classifiers perform comparably to more complex models, information is likely linearly encoded; if random forests or MLPs significantly outperform linear models, information is non-linearly encoded (Belinkov and Glass, 2019).
2.4 Training and Evaluation
We stratify all datasets into train, validation and test splits. We train classifiers using the training set, select hyperparameters using the validation set, and evaluate on the held-out test set using accuracy and macro F1. We solve equation 1 in closed-form to identify $W\in\mathbb{R}^{d\times c}$, tune random forests for number of trees & maximum depth, and learn $W_{1}$ and $W_{2}$ for the MLP via stochastic gradient descent using the AdamW optimizer on cross-entropy loss (Kingma and Ba, 2014) [^4] We use weight decay for regularization for MLP models.
Linear Separability Gap
To quantify whether a layer’s information is linearly encoded or not, we define the linear separability gap as the difference in accuracy between non-linear and linear models. A larger gap indicates the signal is buried in non-linear structure, while a small gap suggests it is directly accessible. In practice, this gap is typically between -0.3 and 0.3, though theoretically it can range from -1 to 1.
Model Parameters Pretraining Data
Encoder-only
BERT Base 110M 12.6B tokens11footnotemark: 1
BERT Large 340M 12.6B tokens11footnotemark: 1
DeBERTa V3 Large 418M 32B tokens11footnotemark: 1
Decoder-only
GPT 2 Small 124M 8B tokens11footnotemark: 1
GPT 2 Large 708M 8B tokens11footnotemark: 1
GPT 2 XL 1.5B 8B tokens11footnotemark: 1
Pythia 6.9b 6900M 300B tokens
Pythia 6.9b Tulu 6900M 300B tokens
OLMo 2 7B 7300M 4T tokens
OLMo 2 7B Instruct 7300M 4T tokens
Gemma 2 2B 2610M 2T tokens
Gemma 2 2B Instruct 2610M 2T tokens
Qwen2.5 1.5B 1540M 18T tokens
Qwen2.5 1.5B Instruct 1540M 18T tokens
Llama 3.1 8B 8000M 15T tokens
Llama 3.1 8B Instruct 8000M 15T tokens
Table 1: Overview of models used in this study. 1Converted from GB to tokens using the approximation that 1GB of data is approximately 200M tokens in English (Chang et al., 2024).
Figure 2: Linguistic accuracy of different classifiers across model layers. The top row shows lexeme prediction accuracy, while the bottom row shows inflection prediction accuracy. Columns represent different classifier types: Linear Regression (left), MLP (middle), and Random Forest (right). Each line represents a different model.
3 Experiments
Using the classifier methodology introduced in § 2, we describe the components of our experimental setup: the dataset, model suite, and procedure for extracting token-level representations.
3.1 Dataset
We use the Universal Dependencies Georgetown University Multilayer (GUM) Corpus in our experiments (Zeldes, 2017). This corpus provides rich annotations for English texts across various genres. We curate a dataset containing 54,816 data points from 8,415 unique sentences, with 8 inflection types, 7,848 unique lemmas, and 11,720 unique word forms.
Each data point corresponds to a target word within its sentence context and includes the full sentence text, the target word’s position in the sentence, its lemma (base form), and its inflection label (grammatical category).
Formally, a data point is a tuple $(S,i,\ell,c)$ where $S=\{w_{0},w_{1},...,w_{n}\}$ is a sentence, $i$ is the index of the target word $w_{i}$ within the sentence, $\ell$ is the lemma of $w_{i}$, and $c$ is the inflection category of $w_{i}$.
The data covers three main syntactic categories: nouns (49.5%), verbs (31.2%), and adjectives (19.4%). The most frequent inflection categories are singular nouns (36.2%), verb base forms (18.4%), and positive adjectives (18.1%), while comparative (0.7%) and superlative (0.5%) forms are less common.
Sentences are relatively concise with an average length of 6.5 words, making them ideal for our analysis as they reduce contextual complexity while providing sufficient context. We split the dataset into training, validation, and test sets with a 70-10-20 ratio using stratified sampling based on inflection labels to ensure balanced representation across all grammatical categories.  [^5] See Appendix B for complete details including tokenization statistics and visualizations
3.2 Models
We study a diverse set of pretrained transformer language models available via the Hugging Face Transformers library (Wolf et al., 2020). Our selection includes different model architectures (encoder-only, decoder-only), sizes and training regimes. Table 1 lists the models used in this study.
3.3 Representation Extraction
For each data point $(S,i,\ell,c)$, we tokenize the sentence $S$ using model-specific tokenizers and process it through the model to obtain hidden states from all layers.  [^6] For target words split into multiple subword tokens, we use the representation of the last subword token to represent the entire word (Devlin et al., 2019). These representations form the basis of our layer-wise analysis of how linguistic information is encoded throughout the model.
4 Results
Figure 3: Classifier selectivity across model layers. Selectivity is defined as the difference between linguistic and control task accuracy, with higher values indicating better generalization rather than memorization. The top row shows lexeme selectivity; the bottom row shows inflection selectivity. Columns represent different classifier types: Linear Regression (left), MLP (middle), and Random Forest (right). Each line represents a different model.
4.1 Lexical Identity
Linear Regression Classifiers
Lemma prediction accuracy using linear regression shows a consistent declining pattern across layer depth for all models, beginning with accuracy between 0.8-1.0 in early layers and declining toward final layers (Figure 2, top left). DeBERTa-v3-Large exhibits the sharpest decline (0.9 $\rightarrow$ 0.3), while Pythia-6.9B maintains the highest sustained performance in later layers. Performance varies significantly across parts of speech, with models achieving highest accuracy on frequent categories like nouns and verbs, while struggling with infrequent categories like pronouns and prepositions due to limited training examples (Table 10).
MLP Classifiers
Two-layer MLP classifiers show higher overall accuracies and less pronounced decline compared to linear regression (Figure 2, top middle). Pythia models show the strongest performance throughout, while GPT-2, BERT and Qwen2.5 variants show the largest improvements from linear to MLP classifiers. The same frequency-dependent performance patterns persist across parts of speech (Table 11).
Random Forest Classifiers
Random forest results for lemma prediction are not reported due to computational constraints arising from thousands of unique lemma classes.
4.2 Inflectional Morphology
Linear Regression Classifiers
Inflectional morphology prediction using linear regression exhibits consistently high performance (0.9-1.0) across all layers and models (Figure 2, bottom left). The accuracy curves show little degradation from early to late layers. Between classifiers trained on different language model representations, there is minimal variance in accuracy. While most morphological forms achieve high accuracy, comparative and superlative adjective forms are most challenging (Table 8).
MLP Classifiers
MLP classifiers show nearly identical patterns to linear regression, maintaining accuracy between 0.9-1.0 across all layers (Figure 2, bottom middle). This highlights the fact that inflectional morphology is encoded very linearly, as additional nonlinear capacity provides no meaningful benefit. The pattern of lower performance on comparative and superlative forms persists with MLP classifiers (Table 9).
Random Forest Classifiers
Random forest classifiers exhibit lower overall accuracy (0.65-0.90) and a declining trend across layers (Figure 2, bottom right), contrasting sharply with the flat, high-performance curves of linear regression and MLP classifiers.
Figure 4: Fraction of maximum components needed to achieve different explained variance thresholds across layers. Each subplot shows a different variance threshold (50%-100%), with lines colored by model. Note that 100% explained variance always requires the full component set. Models fall into two groups: in one, dimensionality declines gradually; in the other, the middle layers require only one component. See additional plots in Figure 8.
Model $d_{\text{model}}$ ID50 ID70 ID90
First Mid Final First Mid Final First Mid Final
BERT-Base 768 123 100 88 244 212 192 461 451 446
BERT-Large 1024 138 105 85 286 226 208 567 527 554
DeBERTa-v3-Large 1024 196 133 29 377 299 113 688 635 423
GPT-2-Small 768 37 1 1 152 1 1 402 1 3
GPT-2-Large 1280 24 1 95 172 1 284 583 1 726
GPT-2-XL 1600 113 1 118 340 1 356 838 1 914
Pythia-6.9B 4096 391 1 96 865 1 517 1952 1 1925
Pythia-6.9B-Tulu 4096 390 1 244 862 1 832 1949 1 2292
OLMo-2-7B 4096 404 310 41 833 896 299 1772 2279 1550
OLMo-2-7B-Instruct 4096 404 358 111 833 974 567 1772 2361 1964
Gemma-2-2B 2304 216 8 11 505 130 70 1129 794 611
Gemma-2-2B-Instruct 2304 222 22 8 520 198 57 1153 899 572
Qwen-2.5-1.5B 1536 184 1 9 399 1 50 835 1 452
Qwen-2.5-1.5B-Instruct 1536 184 1 11 394 1 70 820 1 533
Llama-3.1-8B 4096 373 240 35 789 727 187 1722 2051 1119
Llama-3.1-8B-Instruct 4096 372 215 31 788 664 181 1722 1957 1093
Table 2: Number of principal-component axes required to reach 50% (ID50), 70% (ID70) and 90% (ID90) explained variance in the first, middle and last layers of each model.
5 Analysis
Our experiments reveal clear patterns in how transformer models encode linguistic information. We organize our analysis around six key aspects: distinct encoding patterns, classifier complexity, generalization versus memorization, representation dimensionality, the role of model architectures & training dynamics, and the effects of tokenization.
Distinct Encoding Patterns
We find consistent and distinct encoding patterns for lexical and inflectional information. Lexical information is most accessible in early layers (0-2) and then declines, while inflectional information remains uniformly high across all layers (Figure 2). This separation suggests that lexical identity is captured as an early, surface-level feature, while morphological features remain explicit throughout. This extends the hierarchical view of transformer representations advanced by Tenney et al. (2019) and Jawahar et al. (2019), by showing that surface morphology is not relegated to early layers alone but is preserved as a linearly decodable signal at every depth.
Differences in Classifier Complexity
We find that different types of linguistic information require different levels of classifier complexity to extract (Figure 2), which we quantify using the linear separability gap from § 2. Inflectional features show a small gap across all models, indicating that they remain linearly decodable. Lexical features, by contrast, show an increasingly large gap in later layers, suggesting that lexical features become progressively more non-linearly encoded (Figure 7 in Appendix C).
Generalization vs. Memorization
A key question is whether classifiers detect genuine linguistic structure or simply memorize training examples. We implement control tasks following Hewitt and Liang (2019), where words are randomly but consistently mapped to arbitrary labels. High classifier selectivity - the accuracy difference between linguistic task accuracy and control task accuracy - indicates generalization over memorization.
As shown in Figure 3, inflection prediction shows strong selectivity (0.4-0.6), especially in later layers, suggesting that models develop generalizable abstractions for morphological patterns. In contrast, lemma prediction has near-zero selectivity across all models, layers, and classifier types, suggesting that lexical information is retrieved primarily through memorization. [^7] This is perhaps unsurprising given the nature of the lemma task: it involves thousands of unique classes, many of which appear infrequently in the training data.
This aligns with the notion that morphological rules are systematic and generalizable (e.g., adding -ed to form past tense), while lexical identity is item-specific. It also connects to previous work on the increasing contextualization of representations in later layers of transformer models (Ethayarajh, 2019). Subramani et al. (2022) similarly found that steering vectors become less effective in later layers and that representations reflect semantic similarity less in later layers.
Taken together, these suggest that abstract linguistic properties become increasingly entangled with contextual information, making them increasingly harder to extract with linear methods but still recoverable through non-linear transformations.
Representation Dimensionality
[…]
Models fall into two groups (Table 2). In one group (BERT, DeBERTa, Gemma, Llama 3.1), dimensionality declines gradually across layers. In the other (GPT-2, Qwen2.5, Pythia, OLMo-2), the middle layers require just one component to reach 50-99% variance, before rising again towards the final layers of the model
Importantly, these low-dimensional middle layers are where inflectional features are most linearly separable and where lexical features show the largest non-linear gap (see Figure 7). In other words, middle layers learn to compress representations into a tight subspace that cleanly encodes inflectional morphology but entangles lexemes. Later layers expand dimensionality again to support generation or deeper contextualization while preserving linearly accessible inflectional information.
Model Architectures
Despite differences in architecture (encoder vs. decoder), model size, and training objectives, all models show remarkably consistent patterns: relatively flat, high-accuracy curves for inflection prediction and declining accuracy for lemma prediction in later layers (Figure 2). This holds across causal decoder (GPT-2, Pythia, OLMo-2, Qwen, Gemma, Llama 3.1) and bidirectional encoder models (BERT, DeBERTa).
Surprisingly, instruction-tuned models like Qwen-2.5-Instruct, Llama-3-8B-Instruct, and Pythia-6.9B-Tulu display nearly identical trends to their base counterparts (Appendix CTable 7). This suggests that instruction tuning does not substantially alter how lexical or inflectional features are encoded.
Training Dynamics
We also examined how these patterns develop during pre-training by training classifiers at intermediate model checkpoints for both OLMo-2 and Pythia (Figure 9 and Figure 10). Lexeme classification reaches peak accuracy early in training before slowly declining, while inflectional accuracy remains high and stable across both checkpoints and layer depth. This result suggests that models establish the separation between lexical and inflectional information early and maintain this structure throughout training. Taken together with our model architecture findings, we find that separating lexical and inflectional information facilitates effective next-token prediction.
Effects of Tokenization
Figure 5: Effect of tokenization strategy on analogy completion rank. Each point corresponds to a model (color) and analogy (shape). The x-axis is the rank using whole-word representations. The y-axis is the rank using tokenized representations. Here, rank means the position of the expected word when all vocabulary words are sorted by similarity to the resulting embedding from vector arithmetic; lower is better. Points above the gray $y{=}x$ line mean tokenization hurts performance.
Tokenization is an essential component of language modeling. To test how tokenization influences our findings, we use analogy completion tasks (e.g., man:king::woman:?) and compare two approaches: averaging subtoken embeddings after standard tokenization and summing embeddings from whole-word tokens.
For each approach, we perform vector arithmetic on word representations (e.g., king - man + woman). We measure performance by ranking all vocabulary words by cosine similarity to the resulting representation, and observe how highly the expected word (e.g., queen) ranks, with a lower rank indicating better performance.
Whole-word representations markedly outperform averaged subtokens across all models (Figure 5), implying that linguistic regularities are primarily stored in whole-word embeddings rather than compositionally across subtokens. Despite tokenization effects, our classifier results show consistent patterns across models using different tokenizers (see Table 5), indicating robust encoding of lexical and morphological information.
6 Related Work
Probing for linguistic information
Probing studies use supervised classifiers to predict linguistic properties from model representations (Alain and Bengio, 2017; Adi et al., 2017). Methodological improvements include control tasks to ensure probe selectivity (Hewitt and Liang, 2019). Extensive work has established that transformer models learn hierarchical linguistic structures, with different layers specializing in different information types: lower layers capture surface features and morphology, middle layers encode syntax, and upper layers represent semantics and context (Jawahar et al., 2019; Tenney et al., 2019; Liu et al., 2019; Rogers et al., 2020). Most relevant to our work, Vulić et al. (2020) found that lexical information concentrates in lower layers, while Ethayarajh (2019) showed that representations become increasingly context-specific in higher layers.
Activation steering
Beyond probing, recent work has explored manipulating model behavior by intervening on internal representations. This includes steering vectors (Subramani et al., 2022), inference-time interventions (Li et al., 2023), representation editing (Meng et al., 2022), sparse autoencoders for feature discovery (Bricken et al., 2023), and causal mediation analysis (Vig et al., 2020). These methods complement our classification-based analysis by providing insights into the functional role of different representations. See Appendix A for more detailed discussion of these interpretability approaches.
7 Conclusion
In this work, we examine how 16 transformer models from BERT to Llama 3.1 encode lexical identity and inflectional morphology. To do this, we train linear and non-linear classifiers on layer-wise activations to predict word lemmas and inflectional features, while using control tasks to measure classifier selectivity. Our investigation reveals consistent patterns across all models: lexical information concentrates in early layers and becomes increasingly non-linear in deeper layers, while inflectional information remains linearly accessible throughout the network. Moreover, our selectivity analysis shows that models primarily memorize lexical relationships, but learn generalizable patterns for inflection.
Crucially, these encoding patterns hold across different architectures, model sizes, and training procedures. These findings suggest that this linguistic information is essential for both next token prediction during pretraining and complex instruction-following, demonstrating that despite rapid advances in language technologies, certain aspects of how these models process language remain remarkably consistent.
8 Limitations
Representation Extraction for Decoder Models
Our current approach for extracting word representations from decoder-only models uses the final subword token. This assumption is an intuitive and natural choice, but may not be optimal for all architectures and models. Future work could develop better extraction methods that account for subword tokenization effects and leverage attention patterns to create more accurate word-level representations.
Form and Function in English Inflection
English contains cases where different grammatical functions share the same surface form (e.g., infinitive and non-past verb forms). We do not explicitly look at these cases in our classification/probing experiments, but these ambiguities create opportunities to better examine how models separate form from function.
English-Only Analysis
Our analysis focuses solely on English, limiting generalization to languages with richer morphological systems, where the separation of lexical identity and inflectional morphology may differ in fundamental ways.
Indirect Nature of Classifiers
While our classifier methodology follows established best practices (Hewitt and Liang, 2019; Liu et al., 2019), we only detect correlations in hidden activations, not causal mechanisms. The increasing non-linearity of lexical encoding in deeper layers suggests that linear regression, MLP and random forest classifiers may not fully capture the complete picture of how information is organized.
Appendix A Additional Related Work
A.1 Advanced Probing Methodologies
Beyond standard linear probes, there are many sophisticated approaches to understand model representations. Amnesic probing (Elazar et al., 2021) removes specific information from representations to test whether it’s necessary for downstream tasks. Minimum description length probes (Voita and Titov, 2020) balance probe complexity with performance to avoid overfitting. Causal probing (Geiger et al., 2021) aims to establish causal rather than merely correlational relationships between representations and linguistic properties. Recently, Subramani et al. (2025) find that decoding from activations directly using the Logit Lens can be used to learn confidence estimators for tool-calling agents (nostalgebraist, 2020).
A.2 Model Manipulation and Steering
Steering vectors demonstrate that specific directions in activation space correspond to high-level behavioral changes (Subramani et al., 2022). Building on this, Panickssery et al. (2024) achieves behavioral control by adding activation differences between contrasting examples. Li et al. (2023) introduce inference-time intervention, a method that shifts model activations during inference across limited attention heads to control model behavior. While these methods operate in activation space, task vectors enable arithmetic operations on model capabilities by manipulating weight space (Ilharco et al., 2023).
A.3 Feature Discovery and Mechanistic Interpretability
Recent work has explored using sparse autoencoders to discover latent features (Cunningham et al., 2023; Bricken et al., 2023), providing clearer targets for steering and interpretation. Mechanistic interpretability approaches aim to reverse-engineer the algorithms learned by neural networks (Elhage et al., 2021). Representation editing directly modifies model weights to alter specific factual associations (Meng et al., 2022). These methods complement classification-based approaches by identifying the underlying structure of learned representations and their functional significance.
Appendix B Dataset Statistics
This section provides additional statistics and visualizations for the dataset used in our experiments. Only words containing alphabetic characters and apostrophes were considered. The data covers three main syntactic categories: nouns, verbs, and adjectives, with nouns being most frequent (49.5%). For our experimental setup, we split the dataset into training, validation, and test sets with a 70-10-20 ratio using stratified sampling based on inflection labels.
B.1 Category and Inflection Distribution
Table 3 shows the distribution of word categories in our dataset. Nouns constitute nearly half of all target words, followed by verbs and adjectives.
Count Percentage
Noun 27111 49.5
Verb 17093 31.2
Adjective 10612 19.4
Table 3: Distribution of word categories in the dataset
Table 4 presents the distribution of inflection categories across the dataset. Singular nouns are the most frequent form (36.2%), while adjectival forms like comparative and superlative are relatively rare (0.7% and 0.5% respectively).
Count Percentage
singular 19830 36.2%
base 10076 18.4%
positive 9926 18.1%
plural 7281 13.3%
past 5604 10.2%
3rd_pers 1413 2.6%
comparative 403 0.7%
superlative 283 0.5%
Table 4: Distribution of inflection categories in the dataset
B.2 Sentence Statistics
Sentences in the corpus are relatively concise, with an average length of 6.5 words and a median of 5 words. The minimum number of words is 1 and the maximum is 40. This relatively short average length is ideal for our analysis, as it reduces the complexity of contextual dependencies while still providing sufficient context for meaningful representation.
Figure 6 shows the distribution of sentence lengths.
Figure 6: Distribution of sentence lengths in the dataset
B.3 Tokenization Statistics
Model Tokenizer Type
BERT Base/Large WordPiece
DeBERTa V3 Large SentencePiece
GPT-2 variants BPE
Pythia variants BPE
OLMo 2 variants BPE (tiktoken)
Gemma 2 variants SentencePiece
Qwen 2.5 variants Byte-level BPE
Llama 3.1 variants BPE (tiktoken)
Table 5: Tokenization strategies used by different model families. BPE means byte-pair encoding.
An important consideration for our analysis is how different models tokenize the words in our dataset. Table 6 shows tokenization statistics across the models we analyze. Encoder-only models like BERT and DeBERTa tend to split words into more tokens than decoder-only models like GPT-2 and Qwen2, which may affect how information is encoded across layers.
Avg. tokens Med. tokens Max tokens Percent multitoken
Model per word per word per word
BERT-base 3.11 3.0 8.0 100.0
BERT-large 3.11 3.0 8.0 100.0
DeBERTa-v3-large 3.03 3.0 6.0 100.0
GPT-2-small 1.52 1.0 5.0 42.25
GPT-2-large 1.52 1.0 5.0 42.25
GPT-2-xl 1.52 1.0 5.0 42.25
Pythia-6.9B 1.48 1.0 5.0 39.1
Pythia-6.9B-Tulu 1.48 1.0 5.0 39.1
OLMo2-7B 1.43 1.0 4.0 35.9
OLMo2-7B-Instruct 1.43 1.0 4.0 35.9
Gemma2-2B 2.19 2.0 5.0 100.0
Gemma2-2B-it 2.19 2.0 5.0 100.0
Qwen2.5-1.5B 1.43 1.0 4.0 35.9
Qwen2.5-1.5B-Instruct 1.43 1.0 4.0 35.9
Llama-3.1-8B 2.43 2.0 5.0 100.0
Llama-3.1-8B-Instruct 2.43 2.0 5.0 100.0
Table 6: Tokenization statistics across different models. Open-vocabulary models (GPT-2, Pythia-1.4B, Qwen2) require on average only 1.4–1.5 tokens per word, splitting roughly 36–42% of words into multiple tokens. By contrast, models with closed vocabularies (Gemma-2B, BERT variants, DeBERTa-v3-large) exhibit much finer subword granularity—averaging 2.2–3.1 tokens per word and splitting every word into multiple tokens. Median tokens per word reflect this: 1 for open-vocabulary models, 2–3 for closed-vocabulary models. Maximum tokens per word range from 4 (Qwen2) up to 8 (BERT).
Appendix C Linear Separability Analysis
Lexeme Inflection
Linear MLP Linear MLP
Model Slope R2 Slope R2 Slope R2 Slope R2
BERT-Base $-0.368$ $0.994$ $-0.376$ $0.924$ $0.011$ $0.063$ $0.013$ $0.096$
BERT-Large $-0.414$ $0.977$ $-0.331$ $0.878$ $0.007$ $0.025$ $0.005$ $0.020$
DeBERTa-v3-Large $-0.615$ $0.992$ $-0.527$ $0.935$ $-0.005$ $0.018$ $-0.003$ $0.005$
GPT-2-Small $-0.261$ $0.948$ $-0.049$ $0.006$ $-0.016$ $0.076$ $-0.012$ $0.032$
GPT-2-Large $-0.303$ $0.986$ $-0.389$ $0.860$ $-0.028$ $0.208$ $-0.017$ $0.183$
GPT-2-XL $-0.282$ $0.983$ $-0.367$ $0.868$ $-0.036$ $0.329$ $-0.022$ $0.315$
Pythia-6.9B $-0.213$ $0.992$ $-0.127$ $0.056$ $-0.009$ $0.053$ $-0.007$ $0.028$
Pythia-6.9B-Tulu $-0.195$ $0.988$ $-0.111$ $0.043$ $-0.005$ $0.016$ $-0.002$ $0.003$
OLMo-2-7B $-0.153$ $0.566$ $-0.186$ $0.614$ $0.011$ $0.067$ $0.015$ $0.137$
OLMo-2-7B-Instruct $-0.157$ $0.574$ $-0.178$ $0.615$ $0.009$ $0.051$ $0.013$ $0.106$
Gemma-2-2B $-0.161$ $0.413$ $-0.198$ $0.714$ $0.003$ $0.009$ $0.004$ $0.013$
Gemma-2-2B-Instruct $-0.116$ $0.233$ $-0.207$ $0.505$ $0.016$ $0.176$ $0.014$ $0.138$
Qwen2.5-1.5B $-0.260$ $0.629$ $0.040$ $0.007$ $0.004$ $0.007$ $0.015$ $0.054$
Qwen2.5-1.5B-Instruct $-0.191$ $0.392$ $0.126$ $0.077$ $0.002$ $0.002$ $0.015$ $0.050$
Llama-3.1-8B $-0.109$ $0.188$ $-0.163$ $0.387$ $0.015$ $0.105$ $0.014$ $0.101$
Llama-3.1-8B-Instruct $-0.076$ $0.103$ $-0.121$ $0.258$ $0.018$ $0.153$ $0.016$ $0.138$
Table 7: Regression slopes and R2 values for regressing probe accuracy on layers. Negative slopes indicate decreasing accuracy with layer depth.
Model 3rd person Base Comparative Past Plural Positive Singular Superlative
(n=249) (n=1,833) (n=76) (n=1,003) (n=1,247) (n=1,785) (n=3,587) (n=52)
BERT-Base 0.960 0.965 0.817 0.967 0.983 0.946 0.971 0.759
BERT-Large 0.956 0.964 0.861 0.968 0.982 0.950 0.971 0.768
DeBERTa-v3-Large 0.938 0.974 0.831 0.961 0.986 0.954 0.977 0.706
GPT-2-Small 0.828 0.958 0.840 0.956 0.974 0.941 0.964 0.754
GPT-2-Large 0.812 0.958 0.826 0.951 0.975 0.936 0.967 0.792
GPT-2-XL 0.817 0.959 0.813 0.948 0.977 0.940 0.968 0.788
Pythia-6.9B 0.886 0.972 0.904 0.964 0.989 0.957 0.977 0.907
Pythia-6.9B-Tulu 0.899 0.973 0.909 0.967 0.989 0.956 0.976 0.910
OLMo-2-1124-7B 0.938 0.968 0.902 0.972 0.981 0.923 0.966 0.888
OLMo-2-1124-7B-Instruct 0.927 0.967 0.896 0.971 0.981 0.923 0.965 0.872
Gemma-2-2B 0.901 0.968 0.797 0.969 0.986 0.947 0.974 0.833
Gemma-2-2B-Instruct 0.913 0.966 0.863 0.973 0.988 0.938 0.972 0.872
Qwen2.5-1.5B 0.856 0.950 0.802 0.942 0.972 0.919 0.957 0.688
Qwen2.5-1.5B-Instruct 0.774 0.954 0.647 0.945 0.972 0.921 0.965 0.630
Table 8: Breakdown of inflection classification accuracy by morphological feature for each model, using linear regression classifiers. Inflections are grouped by their morphological features (e.g., Past, Plural, Comparative). For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer. All accuracy values are on a 0–1 scale.
Model 3rd person Base Comparative Past Plural Positive Singular Superlative
(n=249) (n=1,833) (n=76) (n=1,003) (n=1,247) (n=1,785) (n=3,587) (n=52)
BERT-Base 0.973 0.969 0.910 0.972 0.989 0.959 0.974 0.939
BERT-Large 0.967 0.970 0.910 0.973 0.988 0.961 0.975 0.931
DeBERTa-v3-Large 0.954 0.976 0.925 0.966 0.989 0.962 0.979 0.867
GPT-2-Small 0.921 0.963 0.928 0.952 0.972 0.930 0.963 0.870
GPT-2-Large 0.857 0.962 0.872 0.955 0.976 0.942 0.967 0.854
GPT-2-XL 0.921 0.963 0.928 0.952 0.972 0.930 0.963 0.870
Pythia-6.9B 0.932 0.972 0.921 0.961 0.982 0.949 0.971 0.886
Pythia-6.9B-Tulu 0.948 0.974 0.932 0.964 0.983 0.949 0.971 0.897
OLMo-2-1124-7B 0.957 0.968 0.926 0.966 0.989 0.949 0.973 0.905
OLMo-2-1124-7B-Instruct 0.939 0.967 0.903 0.967 0.988 0.949 0.973 0.873
Gemma-2-2B 0.913 0.967 0.863 0.968 0.990 0.950 0.976 0.907
Gemma-2-2B-Instruct 0.930 0.970 0.878 0.975 0.989 0.946 0.974 0.906
Qwen2.5-1.5B 0.882 0.948 0.822 0.943 0.974 0.927 0.957 0.736
Qwen2.5-1.5B-Instruct 0.808 0.953 0.697 0.947 0.974 0.930 0.965 0.682
Table 9: Breakdown of inflection classification accuracy by morphological feature for each model, using MLP classifiers. Inflections are grouped by their morphological features (e.g., Past, Plural, Comparative). For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer. All accuracy values are on a 0–1 scale.
Model Noun Verb Adjective Adverb Pronoun Preposition Conjunction Interjection Other
(n=1,739) (n=641) (n=641) (n=23) (n=1) (n=1) (n=1) (n=1) (n=9)
BERT-Base 0.636 0.737 0.609 0.805 0.292 0.000 0.585 0.000 0.902
BERT-Large 0.684 0.777 0.653 0.826 0.580 0.154 0.662 0.065 0.897
DeBERTa-v3-Large 0.592 0.737 0.585 0.723 0.440 0.077 0.438 0.081 0.866
GPT-2-Small 0.631 0.789 0.612 0.813 0.542 0.000 0.415 0.033 0.896
GPT-2-Large 0.691 0.810 0.688 0.847 0.853 0.174 0.267 0.115 0.912
GPT-2-XL 0.713 0.827 0.708 0.847 0.724 0.222 0.311 0.241 0.899
Pythia-6.9B 0.856 0.926 0.836 0.926 0.938 0.443 0.566 0.488 0.934
Pythia-6.9B-Tulu 0.864 0.930 0.843 0.930 0.923 0.514 0.651 0.476 0.936
OLMo-2-1124-7B 0.798 0.875 0.794 0.913 0.697 0.339 0.363 0.495 0.913
OLMo-2-1124-7B-Instruct 0.798 0.868 0.792 0.902 0.606 0.339 0.331 0.495 0.910
Gemma-2-2B 0.757 0.869 0.736 0.876 0.667 0.179 0.205 0.288 0.891
Gemma-2-2B-Instruct 0.749 0.844 0.742 0.872 0.620 0.137 0.152 0.247 0.912
Qwen2.5-1.5B 0.652 0.801 0.650 0.828 0.526 0.082 0.223 0.068 0.867
Qwen2.5-1.5B-Instruct 0.642 0.800 0.632 0.831 0.544 0.082 0.245 0.068 0.877
Llama-3-8B 0.776 0.882 0.771 0.887 0.831 0.286 0.396 0.321 0.911
Llama-3-8B-Instruct 0.796 0.892 0.788 0.896 0.908 0.300 0.443 0.357 0.917
Table 10: Breakdown of lexeme classification accuracy by Part of Speech (POS) for each model, using linear regression classifiers. Lexemes are grouped by their POS tags (e.g., Noun, Verb, Adjective). For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer. All accuracy values are on a 0–1 scale.
Model Noun Verb Adjective Adverb Pronoun Preposition Conjunction Interjection Other
(n=1,739) (n=641) (n=641) (n=23) (n=1) (n=1) (n=1) (n=1) (n=9)
BERT-Base 0.775 0.831 0.748 0.873 0.458 0.125 0.756 0.267 0.898
BERT-Large 0.813 0.863 0.785 0.884 0.540 0.231 0.725 0.323 0.897
DeBERTa-v3-Large 0.689 0.803 0.682 0.802 0.700 0.115 0.662 0.242 0.861
GPT-2-Small 0.678 0.792 0.665 0.765 0.042 0.000 0.610 0.000 0.830
GPT-2-Large 0.754 0.837 0.755 0.827 0.347 0.188 0.596 0.385 0.871
GPT-2-XL 0.774 0.844 0.771 0.827 0.561 0.232 0.561 0.431 0.860
Pythia-6.9B 0.774 0.856 0.768 0.862 0.554 0.229 0.528 0.310 0.868
Pythia-6.9B-Tulu 0.818 0.880 0.803 0.887 0.554 0.343 0.613 0.381 0.889
OLMo-2-1124-7B 0.818 0.877 0.828 0.896 0.727 0.290 0.734 0.505 0.885
OLMo-2-1124-7B-Instruct 0.822 0.874 0.829 0.897 0.667 0.306 0.750 0.473 0.886
Gemma-2-2B 0.763 0.860 0.763 0.881 0.574 0.125 0.443 0.182 0.880
Gemma-2-2B-Instruct 0.777 0.846 0.785 0.882 0.580 0.137 0.400 0.299 0.875
Qwen2.5-1.5B 0.747 0.838 0.742 0.811 0.228 0.131 0.628 0.164 0.857
Qwen2.5-1.5B-Instruct 0.749 0.840 0.738 0.818 0.211 0.098 0.564 0.123 0.860
Llama-3-8B 0.798 0.879 0.807 0.886 0.800 0.214 0.679 0.393 0.882
Llama-3-8B-Instruct 0.824 0.893 0.826 0.895 0.831 0.257 0.689 0.429 0.887
Table 11: Breakdown of lexeme classification accuracy by Part of Speech (POS) for each model, using Multi-Layer Perceptron (MLP) classifiers. Lexemes are grouped by their POS tags (e.g., Noun, Verb, Adjective). For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer. All accuracy values are on a 0–1 scale.
Linear Separability Gap
(a) Linear separability gap for inflection prediction
(b) Linear separability gap for lemma prediction
Figure 7: Performance advantage of MLP classifiers over linear classifiers (in percentage points) across model layers. The linear separability gap measures how much a non-linear transformation improves classifier performance compared to a simple linear mapping. For inflection prediction, the gap is consistently minimal (mostly within ±$\pm 0.02$ percentage points) and sometimes negative, indicating that inflectional features are primarily encoded in a linear fashion throughout the network. By contrast, the linear separability gap for lemma prediction is relatively large (0.1–0.3 percentage points) and positive across most models
Figure 8: Intrinsic dimensionality curves for all 10 models. Each subplot shows the relationship between the percentage of maximum PCA components (x-axis) and the percentage of explained variance (y-axis) across different layers. The color gradient from purple (early layers, 0%) to yellow (late layers, 100%) indicates the relative layer depth within each model. Models like BERT, Gemma, and Llama show similar compression patterns, while GPT-2 variants, Qwen and Pythia exhibit opposite trends in their middle layers.
Figure 9: Lexeme (top row) and inflection (bottom row) prediction accuracy across normalized layer number for OLMo-2-7B checkpoints at various pretraining steps (5k-734k steps). The full model is 928k steps. Checkpoints are on a color gradient from brightest (earliest) to darkest (latest). Early checkpoints exhibit higher lexeme accuracy than later ones, while inflectional accuracy remains flat across layers and checkpoints.
Figure 10: Lexeme (top row) and inflection (bottom row) prediction accuracy across normalized layer number for Pythia-6.9B checkpoints at various pretraining steps (1-111k steps). The full model is 143k steps. Checkpoints are on a color gradient from brightest (earliest) to darkest (latest). Lexeme accuracy declines both with deeper layers and with more training, whereas inflectional accuracy stays uniformly high across all layers and checkpoints.

### v4

@@
-Model Internal Sleuthing: Finding Lexical Identity and Inflectional Morphology in Modern Language Models
+Echoes of BERT: Do Modern Language Models Rediscover the Classical NLP Pipeline?
 Abstract
-Large transformer-based language models dominate modern NLP, yet our understanding of how they encode linguistic information is rooted in studies of early models like BERT and GPT-2.
-To better understand today’s language models, we investigate how both classical architectures (BERT, DeBERTa, GPT-2) and contemporary large language models (Pythia, OLMo-2, Gemma-2, Qwen2.5, Llama-3.1) represent lexical identity and inflectional morphology.
-We train linear and nonlinear classifiers on layer-wise activations to predict word lemmas and inflectional features.
-We discover that models concentrate lexical information linearly in early layers and increasingly nonlinearly in later layers, while keeping inflectional information uniformly accessible and linearly separable throughout the layers.
-Further analysis reveals that these models encode inflectional morphology through generalizable abstractions, but rely predominantly on memorization to encode lexical identity.
-Remarkably, these patterns emerge across all 16 models we test, despite differences in architecture, size, and training regime (including pretrained and instruction-tuned variants).
-This consistency suggests that, despite substantial advances in LLM technologies, transformer models organize linguistic information in similar ways, indicating that these properties could be fundamental for next token prediction and are learned early during pretraining.
+Large transformer-based language models dominate modern NLP, yet our understanding of how they encode linguistic information relies primarily on studies of early models like BERT and GPT-2.
+Building on prior BERTology work, we analyze 25 models spanning classical architectures (BERT, DeBERTa, GPT-2) to modern large language models (Pythia, OLMo-2, Gemma-2, Qwen2.5, Llama-3.1), probing layer-by-layer representations across eight linguistic tasks in English.
+Consistent with earlier findings, we find that hierarchical organization persists in modern models: early layers capture syntax, middle layers handle semantics and entity-level information, and later layers encode discourse phenomena.
+We dive deeper, conducting an in-depth multilingual analysis of two specific linguistic properties - lexical identity and inflectional morphology - that help disentangle form from meaning.
+We find that lexical information concentrates linearly in early layers but becomes increasingly nonlinear deeper in the network, while inflectional information remains linearly accessible throughout all layers.
+Additional analyses of attention mechanisms, steering vectors, and pretraining checkpoints reveal where this information resides within layers, how it can be functionally manipulated, and how representations evolve during pretraining.
+Taken together, our findings suggest that, even with substantial advances in LLM technologies, transformer models learn to organize linguistic information in similar ways, regardless of model architecture, size, or training regime, indicating that these properties are important for next token prediction.
 Our code is available at https://github.com/ml5885/model_internal_sleuthing
-Model Internal Sleuthing: Finding Lexical Identity and Inflectional Morphology in Modern Language Models
-[…]
 1 Introduction
-Large transformer-based language models (LMs) are widely used for tasks such as text generation, question answering, and code completion (Workshop, 2023; Groeneveld et al., 2024; Grattafiori et al., 2024; Hui et al., 2024).
-However, how these models internally represent linguistic information remains an active research area.
-Prior work suggests a hierarchical organization where different layers specialize in capturing distinct levels of linguistic structure, from surface features to syntax and semantics (Jawahar et al., 2019; Tenney et al., 2019; Rogers et al., 2020).
+Large transformer-based language models (LMs) are widely used for tasks such as text generation, question answering, and code completion (Workshop, 2023; Groeneveld et al., 2024; Llama, 2024; Hui et al., 2024) However, how these models internally represent linguistic information remains an active research area.
+Prior work suggests a hierarchical organization where different layers specialize in capturing distinct levels of linguistic structure, from surface features to syntax and semantics (Jawahar et al., 2019; Tenney et al., 2019a; Rogers et al., 2020).
 But these studies focus only on first-generation LMs such as BERT and GPT-2 (Devlin et al., 2019; Radford et al., 2019).
@@
 We ask: how do modern LMs encode linguistic structure, and how do these representations differ from those in earlier models?
+In particular, do modern LMs rediscover the classical NLP pipeline suggested by prior probing work (Tenney et al., 2019a)?
+In addition to probing for these properties, we perform a focused case study on two linguistic properties: lexical identity and inflectional morphology.
 Consider the words walk, walked, jump, and jumped.
@@
 More broadly, where and how do LMs encode a word’s lemma (its lexical identity) and its grammatical form (inflectional morphology)?
-To answer this, we train classifiers to predict either a word’s lemma or its inflection (e.g., tense, number, degree) from hidden state activations, using annotated sentences from the Universal Dependencies GUM corpus (Zeldes, 2017).
-We evaluate classifiers on 16 language models spanning diverse architectures and sizes.
+To answer these questions, we conduct probing experiments for each of these linguistic properties within model representations (hidden state activations).
+We train classifiers on 25 models spanning different architectures, sizes, and training regimes to determine if they process English linguistic structure in a hierarchical manner.
+For lexical identity and inflectional morphology, we dig deeper, examining six typologically diverse languages: English, Chinese, German, French, Russian, and Turkish.
+We also probe attention heads and pretraining checkpoints, find steering vectors, and analyze intrinsic dimensionality, to show where these two linguistic aspects live within layers, how they evolve during pretraining, and how they can be functionally manipulated.
 We find that:
   1. 1.
-    Lexical information is predominantly encoded in early layers, while inflectional morphology is distributed more uniformly across layers.
+    Modern LMs consistently discover the classical NLP pipeline: early layers process surface and syntactic information, middle layers emphasize semantic and entity-level information, and later layers capture discourse.
   2. 2.
-    These layer-wise distributions hold across all models, despite large-scale changes in architecture, amount of pretraining, and post-training procedures.
+    Lexical information is predominantly encoded in early layers and becomes increasingly non-linear deeper in the network, while inflectional morphology remains linearly accessible across layers and languages.
   3. 3.
-    The effective dimensionality of representations generally declines with layer depth, though in some models the lowest point is in the middle layers - where just one or two dimensions capture nearly all the variance.
-2 Model Internal Sleuthing
+    Lexical identity and inflectional morphology representations emerge early in pretraining, and are robust across architectures, sizes, and languages.
+In particular, we find that inflectional features occupy compact, linearly accessible subspaces in the residual stream, enabling effective steering interventions.
+2 BERTology: Probe Design and Metrics
 Figure 1: Overview of our classifier methodology.
-We extract hidden state activations from each model layer for target words and train two classifiers: one predicting inflection and another predicting lemma.
-We evaluate three classifier architectures (linear regression, MLP, and random forest) for each task to deduce information accessibility and linear separability across the layers of the language model.
-Performance differences between linear and non-linear classifiers reveal the encoding structure of linguistic information.
-To study how language models encode linguistic information, we train simple probabilistic classifiers on the models’ internal representations or activations.
+We extract hidden state activations from each model layer for target words and train classifiers for token, span and pairwise edge predictions (POS, dependencies, constituents, NER, SRL, SPR, coreference, and relations), as well as word-level lemma and inflection prediction.
+We compare linear regression, MLP, and random-forest classifiers, compute selectivity using control labels, and summarize where performance emerges with expected layer and center of gravity.
+We study how language models encode linguistic information by training simple probabilistic classifiers on layer-wise residual-stream activations.
 [^1] We use representations, activations, and neurons interchangeably.
-By observing whether classifiers can recover word-level properties like lemma or inflection from a given representation, we can infer which layers, if any, linguistic information is present, and how directly (e.g., linearly or not).
-High classifier performance at a given layer suggests that the relevant information is encoded and accessible at that layer.
-Additionally, comparing performance across layers tells us how lexical and inflectional information is distributed throughout the model.
-Figure 1 illustrates our approach.
-For each layer of each model, we extract hidden state activations for a target word and use them as input features to train specialized classifiers that predict either inflection or lemma.
-In our experiments, we train three different classifiers: linear regression, two-layer multi-layer perceptron (MLP), and random forest.
-2.1 Linear Regression Classifier
-Consistent with best practices for probing Hewitt and Liang (2019); Liu et al.
+If a classifier recovers a property (e.g., POS, lemma) from a layer $\ell$, we infer that the information is encoded and accessible at that layer; comparing layers shows how it evolves with depth.
+2.1 Probe Design
+We follow the edge probing task design of Tenney et al.
+(2019b).
+For token tasks (e.g., POS), we use the representation of the last subword token for the target word.
+For span tasks (e.g., constituents, NER, SRL, SPR), we represent a span by mean-pooling its subword token representations.
+For pair/edge tasks (e.g., coreference, dependency arcs, relations), we encode a pair $(s_{1},s_{2})$ as $r_{s_{1}};r_{s_{2}};r_{s_{1}}\circ r_{s_{2}};|r_{s_{1}}-r_{s_{2}}|]$, where $r_{s}$ is the token or span representation, $\circ$ is elementwise product, and $|\cdot|$ is absolute value.
+We train linear regression and two-layer multi-layer perceptron (MLP) classifiers with cross-entropy loss on top of representations from each individual layer.
+Unlike Tenney et al.
+(2019a), who uses scalar mixing weights to fuse information across layers, we do not train probes on combinations of multiple layers.
+Our goal is analytical rather than predictive: we care about isolating where in the model each property becomes accessible, rather than achieving the highest possible accuracy.
+Simpler probes trained on individual layers provide a more controlled setting and yield more insightful results about how linguistic information emerges throughout the network.
+See Appendix˜C for details on classifier hyperparameters and training.
+2.1.1 Linear Regression Classifier
+Consistent with best practices for probing Hewitt & Liang (2019); Liu et al.
 (2019), we use simple linear regression classifiers.
@@
 $\hat{Y}_{\text{test}}=X_{\text{test}}W$ (2)
-2.2 Random Forest Classifier
-Random forests provide a complementary approach to linear models by handling complex feature interactions and decision boundaries.
-Unlike linear models that assume information is linearly separable, random forests combine an often large number of decision trees via model averaging to robustly identify patterns (Breiman, 2001).
-Since our classification tasks involve a large number of classes (thousands of unique lemmas and multiple inflection categories), we use a one-vs-all approach.
-For each class $j$, we train a binary random forest classifier that distinguishes class $j$ from all other classes.
-Each binary classifier outputs a probability score $s_{j}(x)\in[0,1]$ representing the confidence that input $x$ belongs to class $j$.
-The final prediction is assigned based on the maximum score across all classes:
-$\hat{y}=\operatorname{argmax}_{j\in\{1,2,...,c\}}s_{j}(x)$ (3)
-where $s_{j}(x)$ is the probability that the trained binary classifier assigns to class $j$ for the input $x$ [^2] We use scikit-learn’s implementation (Pedregosa et al., 2011) with the predict_proba() method to obtain probability scores.
-2.3 MLP Classifier
-To test for more general non-linearity, we train a simple two-layer MLP with ReLU activation, defined as:
+2.1.2 MLP Classifier
+To test for more general nonlinearity, we train a simple two-layer MLP with ReLU activation, defined as:
 $\hat{Y}=\text{softmax}(\text{ReLU}(X_{\text{train}}W_{1})W_{2})$ (3)
@@
 Two-layer MLPs with ReLU activation are universal function approximators, capable of approximating any continuous function to arbitrary precision given sufficient width (Hornik et al., 1989).
-Comparing performance across these three classifier types provides insights into how information is encoded: if linear classifiers perform comparably to more complex models, information is likely linearly encoded; if random forests or MLPs significantly outperform linear models, information is non-linearly encoded (Belinkov and Glass, 2019).
-2.4 Training and Evaluation
+Comparing performance across these classifier types provides insights into how information is encoded: if linear classifiers perform comparably to more complex probes, information is likely linearly encoded; if MLPs significantly outperform linear probes, information is non-linearly encoded (Belinkov & Glass, 2019).
+2.2 Metrics
+We define several metrics for localizing where information emerges with depth and for quantifying nonlinearity: selectivity, the linear separability gap, and two depth statistics inspired by Tenney et al.
+(2019a), expected layer and center of gravity.
+Control tasks and selectivity.
+Probes may simply memorize training data rather than extracting true linguistic information from the representations.
+To measure this, we follow Hewitt & Liang (2019) and train on randomly permuted labels (control tasks).
+Selectivity at layer $\ell$ is the difference between real and control accuracy:
+$\text{Sel}_{\ell}=\text{Acc}^{\text{real}}_{\ell}-\text{Acc}^{\text{control}}_{\ell}$ (4)
+Higher values mean the classifier is extracting true linguistic information rather than memorizing.
+Linear separability gap.
+We quantify nonlinearity at each layer as the difference in accuracy between a non-linear probe and a linear probe:
+$\text{Gap}_{\ell}=\text{Acc}^{\text{nonlin}}_{\ell}-\text{Acc}^{\text{linear}}_{\ell},$ (5)
+where positive values indicate useful information is present but not linearly separable; values near zero suggest linear accessibility.
+The gap lies in [1,1][-1,1] in principle and is typically between 0.3-0.3 and $0.3$ in practice.
+Center of gravity.
+Let $a_{\ell}$ be the test accuracy using layer $\ell$ for $\ell=0,\dots,L$, and let $b_{\ell}=\max_{j\leq\ell}a_{j}$ be the cumulative (best-so-far) curve.
+We weight layers by their consolidation relative to the baseline and take an index-weighted average:
+$w_{\ell}=\frac{b_{\ell}-b_{0}}{\sum_{k=0}^{L}(b_{k}-b_{0})},\quad\text{CenterOfGravity}=\sum_{\ell=0}^{L}\ell\,w_{\ell}.$ (6)
+Higher values indicate that performance is consolidated in higher layers.
+Expected layer.
+To localize where marginal gains arise, we use the nonnegative increments of the cumulative curve and take their index-weighted average:
+$\Delta_{\ell}=\max(b_{\ell}-b_{\ell-1},0)\ ,\quad p_{\ell}=\frac{\Delta_{\ell}}{\sum_{j=1}^{L}\Delta_{j}},\quad\text{ExpectedLayer}=\sum_{\ell=1}^{L}\ell\,p_{\ell}.$ (7)
+Unlike center of gravity (which weights consolidated performance), this emphasizes where useful information first becomes available, highlighting the specific layers at which the model begins to encode properties relevant to the task.
+Model Parameters Pretraining Data Layers
+Encoder-only
+BERT Base 110M 12.6B tokens11footnotemark: 1 12
+BERT Large 340M 12.6B tokens11footnotemark: 1 24
+DeBERTa V3 Large 418M 32B tokens11footnotemark: 1 24
+Decoder-only
+GPT-2-Small 124M 8B tokens11footnotemark: 1 12
+GPT-2-Large 708M 8B tokens11footnotemark: 1 36
+GPT-2-XL 1.5B 8B tokens11footnotemark: 1 48
+Goldfish English 1000mb 124M 200M tokens 12
+Goldfish Chinese 1000mb 124M 200M tokens 12
+Goldfish German 1000mb 124M 200M tokens 12
+Goldfish French 1000mb 124M 200M tokens 12
+Goldfish Russian 1000mb 124M 200M tokens 12
+Goldfish Turkish 1000mb 124M 200M tokens 12
+Pythia-6.9B 6900M 300B tokens 32
+Pythia-6.9B Tulu 6900M 300B tokens 32
+OLMo-2-7B 7300M 4T tokens 32
+OLMo-2-7B-Instruct 7300M 4T tokens 32
+Gemma-2-2B 2610M 2T tokens 26
+Gemma-2-2B-Instruct 2610M 2T tokens 26
+Qwen2.5-1.5B 1540M 18T tokens 28
+Qwen2.5-1.5B-Instruct 1540M 18T tokens 28
+Qwen2.5-7B 7620M 18T tokens 28
+Qwen2.5-7B-Instruct 7620M 18T tokens 28
+Llama-3.1-8B 8000M 15T tokens 32
+Llama-3.1-8B-Instruct 8000M 15T tokens 32
+Encoder-Decoder
+mT5-base 580M 1T tokens 12
+Table 1: Overview of models used in this study
+1Converted from GB to tokens using the approximation that 1GB of data is approximately 200M tokens in English (Chang et al., 2024).
+3 Experiments
+Using the methodology introduced in §˜2, we describe the components of our experimental setup: the datasets, model suite, and procedure for extracting token-level representations.
+3.1 Datasets
+We evaluate the eight tasks used by Tenney et al.
+(2019a): UD English-GUM (POS, dependencies, named entities, coreference, constituents) (Nivre et al., 2016; Zeldes, 2017), Universal Propositions English-EWT (SRL) (Jindal et al., 2022), SPR1 datasets (PropBank and UD-EWT sources; SPR), and SemEval-2010 Task 8 (relations).
+We use the same token/span/edge labeling schemes.
+For our in-depth analysis of lexical identity and inflectional morphology, we use Universal Dependencies corpora across six languages - English, Chinese, German, French, Russian, Turkish (Nivre et al., 2016).
+We select GUM for English (Zeldes, 2017), GSD for Chinese/German/French (McDonald et al., 2013; Guillaume et al., 2019), SynTagRus for Russian (Droganova et al., 2018), and IMST for Turkish (Sulubacak et al., 2016) [^3] See Appendix F for complete details including dataset statistics, tokenization information, and visualizations for all languages
+3.2 Models
+We study a diverse set of pretrained transformer language models spanning different architectures, sizes, and training regimes.
+Table˜1 lists all models used in this study (see Table˜7 for the HuggingFace identifiers).
+For English, we evaluate all models listed in Table˜1 (excluding the non-English Goldfish models).
+For the five non-English languages (Chinese, German, French, Russian, Turkish), we focus on models that have explicit multilingual training: the Goldfish monolingual models trained specifically for each target language (Chang et al., 2024), multilingual Qwen2.5 variants that include these languages in their training data, and the multilingual mT5-base model (Xue et al., 2021).
+This ensures that we evaluate models on languages they were trained on while maintaining sufficient coverage.
+3.3 Representation Extraction
+We tokenize inputs with model-specific tokenizers and run a forward pass to collect residual-stream activations from every layer.
+Token, span, and pair encodings follow §˜2.
+For words split into multiple subwords, we use the last subword’s representation (Devlin et al., 2019).
+Models are used as-is (no fine-tuning), and we report results by layer using these frozen activations.
+4 The Classical NLP Pipeline
+In this section, we present probing results for eight English tasks that span the classical NLP pipeline.
+We report layer-wise accuracy for linear and MLP probes, summarize where each property emerges with Expected Layer and Center of Gravity, and compare models by correlating their layer-by-task accuracy profiles.
+All figures we show here are for three representative models; chosen to cover architecture, size and training regime.
+Full results for all 25 models appear in Appendix˜D.
+4.1 Layer-wise accuracy
+Figure 2: Left: heatmaps of probe accuracy by layer for three models (BERT-Base, Qwen2.5-1.5B, OLMo-2-1124-7B-Instruct).
+The top row shows MLP probe accuracy, the bottom row shows linear probe accuracy.
+Right: Pearson correlations between models are computed by vectorizing each model’s per-layer, per-task accuracy grid.
+The lower triangular matrix shows MLP probe accuracy correlations, the upper triangular matrix shows linear probe accuracy correlations.
+Figure 2 shows accuracy heatmaps for three representative models and correlations between the task accuracy all models.
+In general, we observe high probe accuracy for tasks such as NER and low accuracy for tasks such as semantic relation extraction.
+Mid-level semantic tasks such as SRL and SPR peak in middle layers, while later layers capture discourse-level phenomena like Coreference and Relations.
+To compare results across models, we compute Pearson correlations between the task-by-layer accuracy vectors for each pair of models.
+We find that encoder-only models cluster together, decoder-only models form a separate cluster, and cross-family correlations are weaker.
+With linear probes, model size also separates clusters: small/medium decoder-only models (100M-2B) correlate strongly with each other, while 7B+ models form a distinct high-similarity group.
+4.2 Summary Statistics: Expected Layer and Center of Gravity
+Figure 3: Expected Layer (blue, equation˜7) and Center of Gravity (purple, equation˜6) for the same three models.
+$\tau_{\text{lin}}$ and $\tau_{\text{MLP}}$ give selectivity (real vs.
+control accuracy) for linear and MLP probes.
+Higher $\tau$ indicates probes extract true signal rather than memorizing.
+The summary statistics in Figure 3 demonstrate a more nuanced pattern.
+Consistent with results from Tenney et al.
+(2019a), we find that for early models like BERT-Base and BERT-Large, the hierarchical structure is pronounced: syntax peaks earliest, semantics and entities in the middle, and discourse phenomena latest.
+For larger modern models like Qwen2.5-1.5B and OLMo-2-7B-Instruct, information emerges earlier, indicated by Expected Layer values shifting toward shallower layers.
+These models compress multiple linguistic levels into fewer layers, perhaps reflecting stronger pretraining signals and greater model capacity.
+However, we still observe a hierarchical progression in center-of-gravity metrics, indicating that later layers still play a role in processing linguistic information.
+Selectivity scores confirm that the probe performance reflects real structure rather than memorization, especially for mid- and late-stage tasks.
+4.3 Do Modern LMs Rediscover the Classical NLP Pipeline?
+Overall, our results demonstrate that modern language models consistently rediscover the classical NLP pipeline observed in prior BERTology work.
+Across architectures, training regimes, and model scales, we find that syntactic information is typically represented in earlier layers, semantic information becomes salient in intermediate layers, and discourse-level features are encoded most strongly in later layers.
+However, we also find that model capacity affects where in the network this information emerges.
+Smaller, earlier models such as BERT-Base exhibit the most clearly separated stages, with each linguistic property peaking at a distinct depth.
+As models become larger and more powerful, the peaks for different linguistic properties shift toward earlier layers, indicating that modern models learn rich representations more quickly and need fewer layers to consolidate linguistic knowledge.
+These findings suggest that while the hierarchical organization of linguistic information remains a robust property of language model representations, powerful models build useful representations earlier in the network
+5 Lexical Identity and Inflectional Morphology
+Having established that modern LMs represent the steps of the classical NLP pipeline in a hierarchical manner (§˜4), we now present an in-depth analysis of two token-level properties: lexical identity and inflectional morphology.
+We employ the same probing setup as in §˜4, but expand our analysis to six typologically diverse languages: English, Chinese, German, French, Russian, and Turkish.
+Finally, we examine where and how these properties are encoded in model representations, including which layers make them linearly accessible.
+5.1 Results
+Figure 4: Linguistic accuracy and classifier selectivity across model layers for English.
+The first two columns show lemma (top) and inflection (bottom) prediction accuracy using Linear Regression (left) and MLP (right) classifiers.
+The next two columns show classifier selectivity (difference between linguistic and control task accuracy) for the same tasks and classifiers.
+Higher selectivity indicates better generalization rather than memorization.
+Each line represents a different model.
+Multilingual results are shown in Figure˜5.
+Figure 5: Cross-linguistic patterns in linguistic accuracy and classifier selectivity.
+The first two columns show lemma (top) and inflection (bottom) prediction accuracy using Linear Regression (left) and MLP (right) classifiers.
+The third column shows inflection prediction with Random Forest classifiers (lemma prediction is computationally infeasible due to the large number of classes).
+The rightmost columns show classifier selectivity for lemma (top) and inflection (bottom) tasks.
+Each line represents a different model-language combination.
+We report results for lexical identity and inflectional morphology across classifier types and languages.
+The English set includes 19 models.
+We also evaluate five other languages using 6 multi/monolingual models.
+In addition to MLP and linear regression classifiers, we also use Random Forests for inflection (but not lemma, due to the large number of classes).
+Random Forest results are in Appendix˜I.
+5.1.1 Lemma
+Lemma task accuracy shows a sharp decline in accuracy with depth for linear regression classifiers, starting high (0.8–1.0) in the first layers and dropping in deeper layers (Figure˜4, top left).
+The size of the drop differs by model: DeBERTa-v3-Large falls from about 0.9 to 0.3, while Pythia-6.9B stays stronger later.
+By language, Turkish shows the largest decline (about 0.95 to 0.25), Russian maintains higher accuracy (about 0.6–0.8), and Chinese is in between.
+Two-layer MLPs raise accuracy and soften the decline (Figure˜4, top middle), with Pythia models strongest across layers and GPT-2, BERT, and Qwen2.5 gaining the most from added non-linearity.
+Gains are largest in Turkish, suggesting that in morphologically rich languages lemma information becomes more non-linear deeper in the network.
+Lemma selectivity is near zero across the board, suggesting that the probes mostly memorize.
+5.1.2 Inflectional Morphology
+Inflectional morphology task accuracy remains high (0.9–1.0) across layers and models for linear regression classifiers (Figure˜4, bottom left), with little change with depth across architectures and languages.
+English, German, French, and Russian are in the 0.90–1.0 range, while Turkish is lower (about 0.85–0.92), likely due to many-suffix combinations in agglutinative morphology.
+Comparative and superlative adjectives are consistently hardest.
+MLPs match linear probes for inflection (Figure˜4, bottom middle), suggesting inflection is linearly encoded.
+Random Forests are lower (0.65–0.90) and decline with depth, with Russian and German doing best and Turkish hardest (see Appendix˜I).
+Using control labels, inflection shows clear selectivity (about 0.4–0.6) across models and languages, with the highest values in Russian and German.
+Full curves appear in Figure˜4 and Figure˜5 and tables in Appendix˜J.
+5.1.3 Probe Error Analysis
+We also conduct an error analysis, grouping examples by part of speech and inflection type, and compare performance across frequency of occurrence.
+There is a clear correlation between frequency and accuracy: more frequent lemmas and inflections yield higher accuracy, particularly for lemmas, with rare words consistently the hardest.
+Among inflection types, comparative and superlative degrees, as well as low-count verb forms, exhibit the highest error rates across models.
+These trends are consistent across languages, with Turkish showing the strongest sensitivity to frequency.
+Detailed tables and plots are provided in Appendix˜J.
+5.2 Analysis
+Our experiments demonstrate clear patterns in how models process lexical identity and inflectional morphology.
+Here, we explore further and analyze six key aspects of these representations: cross-linguistic trends, representation structure, attention head outputs vs.
+residual stream, training dynamics, steering experiments, and intrinsic dimensionality.
+Cross-Linguistic Trends
+Across six languages, lemma is most readable early and declines with depth, while inflection stays high across layers (Figure˜5).
+Turkish shows the steepest lemma drop; Russian stays highest overall.
+These trends align with reports that agglutinative systems challenge current LMs.
+Linear Representations
+We compare linear probes with two-layer MLPs using the linear separability gap from §˜2.
+For inflection, gaps are near zero (sometimes slightly positive or negative), indicating that a linear map already captures most of the signal.
+For lemma, gaps are negative in early layers and shrink later, reflecting that regularized linear regression often outperforms a small MLP, especially where lexical information is concentrated and strongly tied to item frequency.
+Intrinsic Dimensionality
+To characterize how representations evolve through depth, we measure intrinsic dimensionality with PCA (Tables˜2 and 11).
+Models fall into two groups: some show gradual dimensionality decline (BERT, DeBERTa, Gemma, Llama 3.1), while others have middle layers where a single component explains most variance (GPT-2, Qwen2.5, Pythia, OLMo-2).
+Some families show gradual decline with depth; others have highly compressed mid-layers.
+Low intrinsic dimensionality has practical implications: we show below that steering vectors can more effectively manipulate model behavior by targeting compressed middle-layer representations.
+Attention vs. Residual Stream
+Comparing averaged attention head outputs to the residual stream, lemma information drops in attention outputs in middle layers while the residual stream stays higher.
+Inflection remains strong in both, with the residual stream slightly better in the middle.
+This supports the view that attention focuses on contextual relations, while MLP blocks store more parametric knowledge.
+See Appendix˜G.
+Training Dynamics
+We evaluate checkpoints for OLMo-2-7B and Pythia-6.9B.
+Three patterns emerge.
+First, the split between lemma and inflection appears early: lemma accuracy peaks at early steps and declines gradually with more training, while inflection accuracy is high and flat from the start.
+Second, the layer at which lemma is most readable shifts slightly toward earlier layers as training proceeds, consistent with increasing contextualization in middle and late layers.
+Third, these trends hold across both families despite different corpora and training scales, pointing to architecture-level behavior rather than a dataset artifact.
+Figures are in Figures˜24 and 25.
+Steering Experiments
+We apply steering vectors computed from class means for inflection categories and test several strengths.
+Steering is effective across most layers and models: mean probability change and flip rates are near 1.0 in many regions, showing that inflectional directions are functionally manipulable.
+The models which are least steerable are the same ones which have a low intrinsic dimensionality in mid-layers, suggesting that compressed representations are harder to adjust with simple vector shifts.
+See Appendix˜H for full details.
+6 Related Work
+Probing for linguistic information
+Probing studies typically use supervised classifiers to predict linguistic properties from model representations (Alain & Bengio, 2017; Adi et al., 2017).
+Extensive work has established that early transformer models (BERT, GPT-2) learn hierarchical linguistic structures, with different layers specializing in different information types: lower layers capture surface features and morphology, middle layers encode syntax, and upper layers represent semantics and context (Jawahar et al., 2019; Tenney et al., 2019a; Rogers et al., 2020).
+More relevant to our work, Vulić et al.
+(2020) found that lexical information concentrates in lower layers, while Ethayarajh (2019) showed that representations become increasingly context-specific in higher layers.
+Activation steering
+Beyond probing, recent work has explored manipulating model behavior by intervening on internal representations.
+This includes steering vectors (Subramani et al., 2022), inference-time interventions (Li et al., 2023), representation editing (Meng et al., 2022), sparse autoencoders for feature discovery (Bricken et al., 2023), and causal mediation analysis (Vig et al., 2020).
+While these methods typically evaluate changes in model outputs, our steering experiments focus on measuring representational changes.
+See Appendix˜B for more detailed discussion.
+7 Conclusion
+In this paper, we examined 25 transformer models and found evidence that modern LMs rediscover the classical NLP pipeline: syntax is most accessible first, then semantic and entity-level information, and finally discourse.
+At the same time, we observe that larger contemporary models shift these peaks toward earlier layers, suggesting that stronger capacity and training lead to useful representations emerging sooner.
+Our in-depth study of lexical identity and inflectional morphology shows a clear pattern.
+Lemma information concentrates early and becomes increasingly non-linear with depth, while inflectional features remain linearly accessible throughout layers and across languages.
+Analyses of attention vs.
+residual streams, pretraining checkpoints, and steering vectors further locate where this information lives, when it appears, and how it can be manipulated.
+These results hold across architectures, sizes, and training regimes, pointing to robust representational regularities that persist despite rapid advances in model scale and training practice.
+8 Reproducibility Statement
+Our GitHub repository is available at https://github.com/ml5885/model_internal_sleuthing.
+It contains code to reproduce dataset construction, probing experiments, and all plots and analyses.
+The main paper specifies the probe design and metrics (§˜2), datasets and model suite (§˜3 and Table˜1), and evaluation summaries for the classical pipeline and for lexical identity and inflectional morphology (§§˜4 and 5).
+The appendix provides complete classifier and training details, dataset statistics, and full-resolution figure grids referenced in the text.
+Together, these materials are intended to enable end-to-end reproduction of our results.
+Appendix A Limitations
+Representation Extraction for Decoder Models
+Our current approach for extracting word representations from decoder-only models uses the final subword token.
+This assumption is an intuitive and natural choice, but may not be optimal for all architectures and models.
+Future work could develop better extraction methods that account for subword tokenization effects and leverage attention patterns to create more accurate word-level representations.
+Form and Function in Inflection
+Some languages contain cases where different grammatical functions share the same surface form (e.g., , infinitive and non-past verb forms in English).
+We do not explicitly examine these cases in our classification experiments, but these ambiguities create opportunities to better examine how models separate form from function across languages.
+Indirect Nature of Classifiers
+While our classifier methodology follows established best practices (Hewitt & Liang, 2019; Liu et al., 2019), we only detect correlations in hidden activations, not causal mechanisms.
+Scope of Steering Experiments
+Our steering vector experiments measure changes in classifier performance rather than downstream model outputs.
+Evaluating effects on actual model generation would require more complex experimental designs to control for confounding factors and ensure that observed changes result from the intended representational modifications rather than other influences.
+Appendix B Additional Related Work
+B.1 Advanced Probing Methodologies
+Beyond standard linear probes, there are many sophisticated approaches to understand model representations.
+Amnesic probing (Elazar et al., 2021) removes specific information from representations to test whether it’s necessary for downstream tasks.
+Minimum description length probes (Voita & Titov, 2020) balance probe complexity with performance to avoid overfitting.
+Causal probing (Geiger et al., 2021) aims to establish causal rather than merely correlational relationships between representations and linguistic properties.
+Recently, Subramani et al.
+(2025) find that decoding from activations directly using the Logit Lens can be used to learn confidence estimators for tool-calling agents (nostalgebraist, 2020).
+B.2 Model Manipulation and Steering
+Steering vectors demonstrate that specific directions in activation space correspond to high-level behavioral changes (Subramani et al., 2022).
+Building on this, Panickssery et al.
+(2024) achieves behavioral control by adding activation differences between contrasting examples.
+Li et al.
+(2023) introduce inference-time intervention, a method that shifts model activations during inference across limited attention heads to control model behavior.
+While these methods operate in activation space, task vectors enable arithmetic operations on model capabilities by manipulating weight space (Ilharco et al., 2023).
+Recent work has also examined how multilingual models like mT5 and ByT5 encode morphological information differently across languages (Dang et al., 2024), finding that tokenization strategies significantly impact morphological representation quality, particularly for morphologically rich languages.
+B.3 Feature Discovery and Mechanistic Interpretability
+Recent work has explored using sparse autoencoders to discover latent features (Cunningham et al., 2023; Bricken et al., 2023), providing clearer targets for steering and interpretation.
+Mechanistic interpretability approaches aim to reverse-engineer the algorithms learned by neural networks (Elhage et al., 2021).
+Representation editing directly modifies model weights to alter specific factual associations (Meng et al., 2022).
+These methods complement classification-based approaches by identifying the underlying structure of learned representations and their functional significance.
+Appendix C Training Details
 We stratify all datasets into train, validation and test splits.
 We train classifiers using the training set, select hyperparameters using the validation set, and evaluate on the held-out test set using accuracy and macro F1.
-We solve equation 1 in closed-form to identify $W\in\mathbb{R}^{d\times c}$, tune random forests for number of trees & maximum depth, and learn $W_{1}$ and $W_{2}$ for the MLP via stochastic gradient descent using the AdamW optimizer on cross-entropy loss (Kingma and Ba, 2014) [^4] We use weight decay for regularization for MLP models.
-Linear Separability Gap
-To quantify whether a layer’s information is linearly encoded or not, we define the linear separability gap as the difference in accuracy between non-linear and linear models.
-A larger gap indicates the signal is buried in non-linear structure, while a small gap suggests it is directly accessible.
-In practice, this gap is typically between -0.3 and 0.3, though theoretically it can range from -1 to 1.
-Model Parameters Pretraining Data
-Encoder-only
-BERT Base 110M 12.6B tokens11footnotemark: 1
-BERT Large 340M 12.6B tokens11footnotemark: 1
-DeBERTa V3 Large 418M 32B tokens11footnotemark: 1
-Decoder-only
-GPT 2 Small 124M 8B tokens11footnotemark: 1
-GPT 2 Large 708M 8B tokens11footnotemark: 1
-GPT 2 XL 1.5B 8B tokens11footnotemark: 1
-Pythia 6.9b 6900M 300B tokens
-Pythia 6.9b Tulu 6900M 300B tokens
-OLMo 2 7B 7300M 4T tokens
-OLMo 2 7B Instruct 7300M 4T tokens
-Gemma 2 2B 2610M 2T tokens
-Gemma 2 2B Instruct 2610M 2T tokens
-Qwen2.5 1.5B 1540M 18T tokens
-Qwen2.5 1.5B Instruct 1540M 18T tokens
-Llama 3.1 8B 8000M 15T tokens
-Llama 3.1 8B Instruct 8000M 15T tokens
-Table 1: Overview of models used in this study. 1Converted from GB to tokens using the approximation that 1GB of data is approximately 200M tokens in English (Chang et al., 2024).
-Figure 2: Linguistic accuracy of different classifiers across model layers.
-The top row shows lexeme prediction accuracy, while the bottom row shows inflection prediction accuracy.
-Columns represent different classifier types: Linear Regression (left), MLP (middle), and Random Forest (right).
-Each line represents a different model.
-3 Experiments
-Using the classifier methodology introduced in § 2, we describe the components of our experimental setup: the dataset, model suite, and procedure for extracting token-level representations.
-3.1 Dataset
-We use the Universal Dependencies Georgetown University Multilayer (GUM) Corpus in our experiments (Zeldes, 2017).
-This corpus provides rich annotations for English texts across various genres.
-We curate a dataset containing 54,816 data points from 8,415 unique sentences, with 8 inflection types, 7,848 unique lemmas, and 11,720 unique word forms.
-Each data point corresponds to a target word within its sentence context and includes the full sentence text, the target word’s position in the sentence, its lemma (base form), and its inflection label (grammatical category).
-Formally, a data point is a tuple $(S,i,\ell,c)$ where $S=\{w_{0},w_{1},...,w_{n}\}$ is a sentence, $i$ is the index of the target word $w_{i}$ within the sentence, $\ell$ is the lemma of $w_{i}$, and $c$ is the inflection category of $w_{i}$.
-The data covers three main syntactic categories: nouns (49.5%), verbs (31.2%), and adjectives (19.4%).
-The most frequent inflection categories are singular nouns (36.2%), verb base forms (18.4%), and positive adjectives (18.1%), while comparative (0.7%) and superlative (0.5%) forms are less common.
-Sentences are relatively concise with an average length of 6.5 words, making them ideal for our analysis as they reduce contextual complexity while providing sufficient context.
-We split the dataset into training, validation, and test sets with a 70-10-20 ratio using stratified sampling based on inflection labels to ensure balanced representation across all grammatical categories.
-[^5] See Appendix B for complete details including tokenization statistics and visualizations
-3.2 Models
-We study a diverse set of pretrained transformer language models available via the Hugging Face Transformers library (Wolf et al., 2020).
-Our selection includes different model architectures (encoder-only, decoder-only), sizes and training regimes.
-Table 1 lists the models used in this study.
-3.3 Representation Extraction
-For each data point $(S,i,\ell,c)$, we tokenize the sentence $S$ using model-specific tokenizers and process it through the model to obtain hidden states from all layers.
-[^6] For target words split into multiple subword tokens, we use the representation of the last subword token to represent the entire word (Devlin et al., 2019).
-These representations form the basis of our layer-wise analysis of how linguistic information is encoded throughout the model.
-4 Results
-Figure 3: Classifier selectivity across model layers.
-Selectivity is defined as the difference between linguistic and control task accuracy, with higher values indicating better generalization rather than memorization.
-The top row shows lexeme selectivity; the bottom row shows inflection selectivity.
-Columns represent different classifier types: Linear Regression (left), MLP (middle), and Random Forest (right).
-Each line represents a different model.
-4.1 Lexical Identity
-Linear Regression Classifiers
-Lemma prediction accuracy using linear regression shows a consistent declining pattern across layer depth for all models, beginning with accuracy between 0.8-1.0 in early layers and declining toward final layers (Figure 2, top left).
-DeBERTa-v3-Large exhibits the sharpest decline (0.9 $\rightarrow$ 0.3), while Pythia-6.9B maintains the highest sustained performance in later layers.
-Performance varies significantly across parts of speech, with models achieving highest accuracy on frequent categories like nouns and verbs, while struggling with infrequent categories like pronouns and prepositions due to limited training examples (Table 10).
-MLP Classifiers
-Two-layer MLP classifiers show higher overall accuracies and less pronounced decline compared to linear regression (Figure 2, top middle).
-Pythia models show the strongest performance throughout, while GPT-2, BERT and Qwen2.5 variants show the largest improvements from linear to MLP classifiers.
-The same frequency-dependent performance patterns persist across parts of speech (Table 11).
-Random Forest Classifiers
-Random forest results for lemma prediction are not reported due to computational constraints arising from thousands of unique lemma classes.
-4.2 Inflectional Morphology
-Linear Regression Classifiers
-Inflectional morphology prediction using linear regression exhibits consistently high performance (0.9-1.0) across all layers and models (Figure 2, bottom left).
-The accuracy curves show little degradation from early to late layers.
-Between classifiers trained on different language model representations, there is minimal variance in accuracy.
-While most morphological forms achieve high accuracy, comparative and superlative adjective forms are most challenging (Table 8).
-MLP Classifiers
-MLP classifiers show nearly identical patterns to linear regression, maintaining accuracy between 0.9-1.0 across all layers (Figure 2, bottom middle).
-This highlights the fact that inflectional morphology is encoded very linearly, as additional nonlinear capacity provides no meaningful benefit.
-The pattern of lower performance on comparative and superlative forms persists with MLP classifiers (Table 9).
-Random Forest Classifiers
-Random forest classifiers exhibit lower overall accuracy (0.65-0.90) and a declining trend across layers (Figure 2, bottom right), contrasting sharply with the flat, high-performance curves of linear regression and MLP classifiers.
-Figure 4: Fraction of maximum components needed to achieve different explained variance thresholds across layers.
-Each subplot shows a different variance threshold (50%-100%), with lines colored by model.
-Note that 100% explained variance always requires the full component set.
-Models fall into two groups: in one, dimensionality declines gradually; in the other, the middle layers require only one component.
-See additional plots in Figure 8.
+For linear regression, we use ridge regularization with $\lambda=0.01$.
+For random forests, we tune the number of trees (50, 100, 200) and maximum depth (5, 10, 20, None) using grid search.
+For MLPs, we use a hidden layer size of 64, learning rate of 0.001, weight decay of 0.01, and train for up to 100 epochs with early stopping based on validation loss.
+We solve equation˜1 in closed-form to identify $W\in\mathbb{R}^{d\times c}$, and learn $W_{1}$ and $W_{2}$ for the MLP via stochastic gradient descent using the AdamW optimizer on cross-entropy loss (Kingma & Ba, 2014).
+[^4] We use weight decay for regularization for MLP models.roximators, capable of approximating any continuous function to arbitrary precision given sufficient width (Hornik et al., 1989).
+Appendix D Full Results for the Classical NLP Pipeline
+Figure 6: Full heatmaps for MLP probe accuracy across all tasks, models, and layers.
+Rows show tasks; columns show models; each cell shows accuracy by layer depth.
+Early layers favor syntax, middle layers semantics, late layers discourse.
+Figure 7: Full heatmaps for linear probe accuracy across all tasks, models, and layers. Trends mirror the MLP version but with stronger model-size effects in deeper layers.
+Figure 8: Full heatmaps for MLP probe selectivity (real vs. control task accuracy). Higher values indicate true signal rather than memorization.
+Figure 9: Full heatmaps for linear probe selectivity (real vs. control task accuracy). Selectivity is generally higher for semantic and discourse tasks.
+Figure 10: Full Expected Layer and Center of Gravity plots with $\tau_{\text{lin}}$ and $\tau_{\text{MLP}}$ selectivity scores for all models.
+We observe early peaks for syntax, middle peaks for semantics, and late peaks for discourse across architectures and sizes.
+Appendix E Additional Analysis
+Figure 11: Intrinsic dimensionality across layers.
+Lines show fraction of PCA components needed to reach variance thresholds (50–100%).
+Models with strong mid-layer compression (few components for high variance) align with the inflection-is-linear, lemma-is-nonlinear split.
+Full curves per model appear in Figure˜12.
+E.1 Intrinsic Dimensionality Results
 Model $d_{\text{model}}$ ID50 ID70 ID90
@@
 Table 2: Number of principal-component axes required to reach 50% (ID50), 70% (ID70) and 90% (ID90) explained variance in the first, middle and last layers of each model.
-5 Analysis
-Our experiments reveal clear patterns in how transformer models encode linguistic information.
-We organize our analysis around six key aspects: distinct encoding patterns, classifier complexity, generalization versus memorization, representation dimensionality, the role of model architectures & training dynamics, and the effects of tokenization.
-Distinct Encoding Patterns
-We find consistent and distinct encoding patterns for lexical and inflectional information.
-Lexical information is most accessible in early layers (0-2) and then declines, while inflectional information remains uniformly high across all layers (Figure 2).
-This separation suggests that lexical identity is captured as an early, surface-level feature, while morphological features remain explicit throughout.
-This extends the hierarchical view of transformer representations advanced by Tenney et al.
-(2019) and Jawahar et al.
-(2019), by showing that surface morphology is not relegated to early layers alone but is preserved as a linearly decodable signal at every depth.
-Differences in Classifier Complexity
-We find that different types of linguistic information require different levels of classifier complexity to extract (Figure 2), which we quantify using the linear separability gap from § 2.
-Inflectional features show a small gap across all models, indicating that they remain linearly decodable.
-Lexical features, by contrast, show an increasingly large gap in later layers, suggesting that lexical features become progressively more non-linearly encoded (Figure 7 in Appendix C).
-Generalization vs. Memorization
-A key question is whether classifiers detect genuine linguistic structure or simply memorize training examples.
-We implement control tasks following Hewitt and Liang (2019), where words are randomly but consistently mapped to arbitrary labels.
-High classifier selectivity - the accuracy difference between linguistic task accuracy and control task accuracy - indicates generalization over memorization.
-As shown in Figure 3, inflection prediction shows strong selectivity (0.4-0.6), especially in later layers, suggesting that models develop generalizable abstractions for morphological patterns.
-In contrast, lemma prediction has near-zero selectivity across all models, layers, and classifier types, suggesting that lexical information is retrieved primarily through memorization.
-[^7] This is perhaps unsurprising given the nature of the lemma task: it involves thousands of unique classes, many of which appear infrequently in the training data.
-This aligns with the notion that morphological rules are systematic and generalizable (e.g., adding -ed to form past tense), while lexical identity is item-specific.
-It also connects to previous work on the increasing contextualization of representations in later layers of transformer models (Ethayarajh, 2019).
-Subramani et al.
-(2022) similarly found that steering vectors become less effective in later layers and that representations reflect semantic similarity less in later layers.
-Taken together, these suggest that abstract linguistic properties become increasingly entangled with contextual information, making them increasingly harder to extract with linear methods but still recoverable through non-linear transformations.
-Representation Dimensionality
-[…]
-Models fall into two groups (Table 2).
-In one group (BERT, DeBERTa, Gemma, Llama 3.1), dimensionality declines gradually across layers.
-In the other (GPT-2, Qwen2.5, Pythia, OLMo-2), the middle layers require just one component to reach 50-99% variance, before rising again towards the final layers of the model
-Importantly, these low-dimensional middle layers are where inflectional features are most linearly separable and where lexical features show the largest non-linear gap (see Figure 7).
-In other words, middle layers learn to compress representations into a tight subspace that cleanly encodes inflectional morphology but entangles lexemes.
-Later layers expand dimensionality again to support generation or deeper contextualization while preserving linearly accessible inflectional information.
-Model Architectures
-Despite differences in architecture (encoder vs.
-decoder), model size, and training objectives, all models show remarkably consistent patterns: relatively flat, high-accuracy curves for inflection prediction and declining accuracy for lemma prediction in later layers (Figure 2).
-This holds across causal decoder (GPT-2, Pythia, OLMo-2, Qwen, Gemma, Llama 3.1) and bidirectional encoder models (BERT, DeBERTa).
-Surprisingly, instruction-tuned models like Qwen-2.5-Instruct, Llama-3-8B-Instruct, and Pythia-6.9B-Tulu display nearly identical trends to their base counterparts (Appendix CTable 7).
-This suggests that instruction tuning does not substantially alter how lexical or inflectional features are encoded.
-Training Dynamics
-We also examined how these patterns develop during pre-training by training classifiers at intermediate model checkpoints for both OLMo-2 and Pythia (Figure 9 and Figure 10).
-Lexeme classification reaches peak accuracy early in training before slowly declining, while inflectional accuracy remains high and stable across both checkpoints and layer depth.
-This result suggests that models establish the separation between lexical and inflectional information early and maintain this structure throughout training.
-Taken together with our model architecture findings, we find that separating lexical and inflectional information facilitates effective next-token prediction.
-Effects of Tokenization
-Figure 5: Effect of tokenization strategy on analogy completion rank.
-Each point corresponds to a model (color) and analogy (shape).
-The x-axis is the rank using whole-word representations.
-The y-axis is the rank using tokenized representations.
-Here, rank means the position of the expected word when all vocabulary words are sorted by similarity to the resulting embedding from vector arithmetic; lower is better.
-Points above the gray $y{=}x$ line mean tokenization hurts performance.
-Tokenization is an essential component of language modeling.
-To test how tokenization influences our findings, we use analogy completion tasks (e.g., man:king::woman:?) and compare two approaches: averaging subtoken embeddings after standard tokenization and summing embeddings from whole-word tokens.
-For each approach, we perform vector arithmetic on word representations (e.g., king - man + woman).
-We measure performance by ranking all vocabulary words by cosine similarity to the resulting representation, and observe how highly the expected word (e.g., queen) ranks, with a lower rank indicating better performance.
-Whole-word representations markedly outperform averaged subtokens across all models (Figure 5), implying that linguistic regularities are primarily stored in whole-word embeddings rather than compositionally across subtokens.
-Despite tokenization effects, our classifier results show consistent patterns across models using different tokenizers (see Table 5), indicating robust encoding of lexical and morphological information.
-6 Related Work
-Probing for linguistic information
-Probing studies use supervised classifiers to predict linguistic properties from model representations (Alain and Bengio, 2017; Adi et al., 2017).
-Methodological improvements include control tasks to ensure probe selectivity (Hewitt and Liang, 2019).
-Extensive work has established that transformer models learn hierarchical linguistic structures, with different layers specializing in different information types: lower layers capture surface features and morphology, middle layers encode syntax, and upper layers represent semantics and context (Jawahar et al., 2019; Tenney et al., 2019; Liu et al., 2019; Rogers et al., 2020).
-Most relevant to our work, Vulić et al.
-(2020) found that lexical information concentrates in lower layers, while Ethayarajh (2019) showed that representations become increasingly context-specific in higher layers.
-Activation steering
-Beyond probing, recent work has explored manipulating model behavior by intervening on internal representations.
-This includes steering vectors (Subramani et al., 2022), inference-time interventions (Li et al., 2023), representation editing (Meng et al., 2022), sparse autoencoders for feature discovery (Bricken et al., 2023), and causal mediation analysis (Vig et al., 2020).
-These methods complement our classification-based analysis by providing insights into the functional role of different representations.
-See Appendix A for more detailed discussion of these interpretability approaches.
-7 Conclusion
-In this work, we examine how 16 transformer models from BERT to Llama 3.1 encode lexical identity and inflectional morphology.
-To do this, we train linear and non-linear classifiers on layer-wise activations to predict word lemmas and inflectional features, while using control tasks to measure classifier selectivity.
-Our investigation reveals consistent patterns across all models: lexical information concentrates in early layers and becomes increasingly non-linear in deeper layers, while inflectional information remains linearly accessible throughout the network.
-Moreover, our selectivity analysis shows that models primarily memorize lexical relationships, but learn generalizable patterns for inflection.
-Crucially, these encoding patterns hold across different architectures, model sizes, and training procedures.
-These findings suggest that this linguistic information is essential for both next token prediction during pretraining and complex instruction-following, demonstrating that despite rapid advances in language technologies, certain aspects of how these models process language remain remarkably consistent.
-8 Limitations
-Representation Extraction for Decoder Models
-Our current approach for extracting word representations from decoder-only models uses the final subword token.
-This assumption is an intuitive and natural choice, but may not be optimal for all architectures and models.
-Future work could develop better extraction methods that account for subword tokenization effects and leverage attention patterns to create more accurate word-level representations.
-Form and Function in English Inflection
-English contains cases where different grammatical functions share the same surface form (e.g., infinitive and non-past verb forms).
-We do not explicitly look at these cases in our classification/probing experiments, but these ambiguities create opportunities to better examine how models separate form from function.
-English-Only Analysis
-Our analysis focuses solely on English, limiting generalization to languages with richer morphological systems, where the separation of lexical identity and inflectional morphology may differ in fundamental ways.
-Indirect Nature of Classifiers
-While our classifier methodology follows established best practices (Hewitt and Liang, 2019; Liu et al., 2019), we only detect correlations in hidden activations, not causal mechanisms.
-The increasing non-linearity of lexical encoding in deeper layers suggests that linear regression, MLP and random forest classifiers may not fully capture the complete picture of how information is organized.
-Appendix A Additional Related Work
-A.1 Advanced Probing Methodologies
-Beyond standard linear probes, there are many sophisticated approaches to understand model representations.
-Amnesic probing (Elazar et al., 2021) removes specific information from representations to test whether it’s necessary for downstream tasks.
-Minimum description length probes (Voita and Titov, 2020) balance probe complexity with performance to avoid overfitting.
-Causal probing (Geiger et al., 2021) aims to establish causal rather than merely correlational relationships between representations and linguistic properties.
-Recently, Subramani et al.
-(2025) find that decoding from activations directly using the Logit Lens can be used to learn confidence estimators for tool-calling agents (nostalgebraist, 2020).
-A.2 Model Manipulation and Steering
-Steering vectors demonstrate that specific directions in activation space correspond to high-level behavioral changes (Subramani et al., 2022).
-Building on this, Panickssery et al.
-(2024) achieves behavioral control by adding activation differences between contrasting examples.
-Li et al.
-(2023) introduce inference-time intervention, a method that shifts model activations during inference across limited attention heads to control model behavior.
-While these methods operate in activation space, task vectors enable arithmetic operations on model capabilities by manipulating weight space (Ilharco et al., 2023).
-A.3 Feature Discovery and Mechanistic Interpretability
-Recent work has explored using sparse autoencoders to discover latent features (Cunningham et al., 2023; Bricken et al., 2023), providing clearer targets for steering and interpretation.
-Mechanistic interpretability approaches aim to reverse-engineer the algorithms learned by neural networks (Elhage et al., 2021).
-Representation editing directly modifies model weights to alter specific factual associations (Meng et al., 2022).
-These methods complement classification-based approaches by identifying the underlying structure of learned representations and their functional significance.
-Appendix B Dataset Statistics
-This section provides additional statistics and visualizations for the dataset used in our experiments.
+Figure 12: Intrinsic dimensionality curves for all models for English.
+Each subplot shows the relationship between the percentage of maximum PCA components (x-axis) and the percentage of explained variance (y-axis) across different layers.
+The color gradient from purple (early layers, 0%) to yellow (late layers, 100%) indicates the relative layer depth within each model.
+Models like BERT, Gemma, and Llama show similar compression patterns, while GPT-2 variants, Qwen and Pythia exhibit opposite trends in their middle layers.
+Appendix F Dataset Statistics
+This section provides statistics and visualizations for the datasets and models used in our experiments across all six languages.
 Only words containing alphabetic characters and apostrophes were considered.
-The data covers three main syntactic categories: nouns, verbs, and adjectives, with nouns being most frequent (49.5%).
-For our experimental setup, we split the dataset into training, validation, and test sets with a 70-10-20 ratio using stratified sampling based on inflection labels.
-B.1 Category and Inflection Distribution
-Table 3 shows the distribution of word categories in our dataset. Nouns constitute nearly half of all target words, followed by verbs and adjectives.
+Language Total Words Unique Lemmas Unique Forms Inflection Types Sentences Avg. Length
+English 54,816 7,848 11,720 8 8,415 6.5
+Chinese 44,166 11,184 11,237 4 7,892 5.8
+German 84,710 24,140 31,890 9 9,234 7.3
+French 115,847 13,804 24,485 6 8,765 6.6
+Russian 193,320 20,943 59,830 8 10,234 7.1
+Turkish 20,881 3,776 11,680 7 6,789 6.4
+Table 3: Dataset statistics across all six languages.
+Russian has the largest dataset and the highest number of unique forms, reflecting its rich inflectional morphology.
+Turkish has the fewest total words and lemmas, while Chinese has the fewest inflection types.
+F.1 English Dataset Details
+For the English GUM corpus specifically, the data covers three main syntactic categories: nouns (49.5%), verbs (31.2%), and adjectives (19.4%).
+Table˜4(a) shows the distribution of word categories in the English dataset, and Table˜4(b) presents the distribution of inflection categories.
 Count Percentage
@@
 Adjective 10612 19.4
-Table 3: Distribution of word categories in the dataset
-Table 4 presents the distribution of inflection categories across the dataset.
-Singular nouns are the most frequent form (36.2%), while adjectival forms like comparative and superlative are relatively rare (0.7% and 0.5% respectively).
+(a) Distribution of word categories
 Count Percentage
@@
 Past 5604 10.2%
-3rd_pers 1413 2.6%
+3rd Person 1413 2.6%
 Comparative 403 0.7%
 Superlative 283 0.5%
-Table 4: Distribution of inflection categories in the dataset
-B.2 Sentence Statistics
-Sentences in the corpus are relatively concise, with an average length of 6.5 words and a median of 5 words.
-The minimum number of words is 1 and the maximum is 40.
-This relatively short average length is ideal for our analysis, as it reduces the complexity of contextual dependencies while still providing sufficient context for meaningful representation.
-Figure 6 shows the distribution of sentence lengths.
-Figure 6: Distribution of sentence lengths in the dataset
-B.3 Tokenization Statistics
+(b) Distribution of inflection categories
+Value
+Average Words 6.5
+Median Words 5
+Minimum Words 1
+Maximum Words 40
+(c) Sentence length statistics
+Table 4: Distribution statistics for the English dataset
+F.2 Tokenization Statistics
 Model Tokenizer Type
@@
 An important consideration for our analysis is how different models tokenize the words in our dataset.
-Table 6 shows tokenization statistics across the models we analyze.
+Table˜6 shows tokenization statistics across the models we analyze.
 Encoder-only models like BERT and DeBERTa tend to split words into more tokens than decoder-only models like GPT-2 and Qwen2, which may affect how information is encoded across layers.
-Avg. tokens Med. tokens Max tokens Percent multitoken
-Model per word per word per word
-BERT-base 3.11 3.0 8.0 100.0
-BERT-large 3.11 3.0 8.0 100.0
-DeBERTa-v3-large 3.03 3.0 6.0 100.0
-GPT-2-small 1.52 1.0 5.0 42.25
-GPT-2-large 1.52 1.0 5.0 42.25
-GPT-2-xl 1.52 1.0 5.0 42.25
-Pythia-6.9B 1.48 1.0 5.0 39.1
-Pythia-6.9B-Tulu 1.48 1.0 5.0 39.1
-OLMo2-7B 1.43 1.0 4.0 35.9
-OLMo2-7B-Instruct 1.43 1.0 4.0 35.9
-Gemma2-2B 2.19 2.0 5.0 100.0
-Gemma2-2B-it 2.19 2.0 5.0 100.0
-Qwen2.5-1.5B 1.43 1.0 4.0 35.9
-Qwen2.5-1.5B-Instruct 1.43 1.0 4.0 35.9
-Llama-3.1-8B 2.43 2.0 5.0 100.0
-Llama-3.1-8B-Instruct 2.43 2.0 5.0 100.0
-Table 6: Tokenization statistics across different models.
-Open-vocabulary models (GPT-2, Pythia-1.4B, Qwen2) require on average only 1.4–1.5 tokens per word, splitting roughly 36–42% of words into multiple tokens.
-By contrast, models with closed vocabularies (Gemma-2B, BERT variants, DeBERTa-v3-large) exhibit much finer subword granularity—averaging 2.2–3.1 tokens per word and splitting every word into multiple tokens.
-Median tokens per word reflect this: 1 for open-vocabulary models, 2–3 for closed-vocabulary models.
-Maximum tokens per word range from 4 (Qwen2) up to 8 (BERT).
-Appendix C Linear Separability Analysis
-Lexeme Inflection
-Linear MLP Linear MLP
-Model Slope R2 Slope R2 Slope R2 Slope R2
-BERT-Base $-0.368$ $0.994$ $-0.376$ $0.924$ $0.011$ $0.063$ $0.013$ $0.096$
-BERT-Large $-0.414$ $0.977$ $-0.331$ $0.878$ $0.007$ $0.025$ $0.005$ $0.020$
-DeBERTa-v3-Large $-0.615$ $0.992$ $-0.527$ $0.935$ $-0.005$ $0.018$ $-0.003$ $0.005$
-GPT-2-Small $-0.261$ $0.948$ $-0.049$ $0.006$ $-0.016$ $0.076$ $-0.012$ $0.032$
-GPT-2-Large $-0.303$ $0.986$ $-0.389$ $0.860$ $-0.028$ $0.208$ $-0.017$ $0.183$
-GPT-2-XL $-0.282$ $0.983$ $-0.367$ $0.868$ $-0.036$ $0.329$ $-0.022$ $0.315$
-Pythia-6.9B $-0.213$ $0.992$ $-0.127$ $0.056$ $-0.009$ $0.053$ $-0.007$ $0.028$
-Pythia-6.9B-Tulu $-0.195$ $0.988$ $-0.111$ $0.043$ $-0.005$ $0.016$ $-0.002$ $0.003$
-OLMo-2-7B $-0.153$ $0.566$ $-0.186$ $0.614$ $0.011$ $0.067$ $0.015$ $0.137$
-OLMo-2-7B-Instruct $-0.157$ $0.574$ $-0.178$ $0.615$ $0.009$ $0.051$ $0.013$ $0.106$
-Gemma-2-2B $-0.161$ $0.413$ $-0.198$ $0.714$ $0.003$ $0.009$ $0.004$ $0.013$
-Gemma-2-2B-Instruct $-0.116$ $0.233$ $-0.207$ $0.505$ $0.016$ $0.176$ $0.014$ $0.138$
-Qwen2.5-1.5B $-0.260$ $0.629$ $0.040$ $0.007$ $0.004$ $0.007$ $0.015$ $0.054$
-Qwen2.5-1.5B-Instruct $-0.191$ $0.392$ $0.126$ $0.077$ $0.002$ $0.002$ $0.015$ $0.050$
-Llama-3.1-8B $-0.109$ $0.188$ $-0.163$ $0.387$ $0.015$ $0.105$ $0.014$ $0.101$
-Llama-3.1-8B-Instruct $-0.076$ $0.103$ $-0.121$ $0.258$ $0.018$ $0.153$ $0.016$ $0.138$
-Table 7: Regression slopes and R2 values for regressing probe accuracy on layers. Negative slopes indicate decreasing accuracy with layer depth.
+Model Avg. tokens Med. tokens Max tokens Percent multitoken
+per word per word per word
+BERT variants 1.11 1.0 6.0 6.95
+DeBERTa-v3-large 1.03 1.0 4.0 2.2
+GPT-2 variants 1.52 1.0 5.0 42.25
+Pythia-6.9B variants 1.48 1.0 5.0 39.1
+OLMo2-7B variants 1.43 1.0 4.0 35.9
+Gemma2-2B variants 1.19 1.0 4.0 16.55
+Qwen2.5-1.5B variants 1.43 1.0 4.0 35.9
+Llama-3.1-8B variants 1.43 1.0 4.0 35.85
+Table 6: Tokenization statistics across different models (English only).
+Most models have an average of 1.0-1.5 tokens per word and a median of 1, indicating that most words are tokenized as a single unit.
+However, there is variation in the proportion of words split into multiple tokens.
+Decoder-only models (e.g., , GPT-2, Pythia, Qwen2, LLaMA) split 35-42% of words, while BERT and DeBERTa variants split fewer words (2-7%).
+Maximum tokens per word range from 4 to 6 across all models.
+F.3 Effects of Tokenization
+Figure 13: Effect of tokenization strategy on analogy completion rank.
+Each point corresponds to a model (color) and analogy (shape).
+The x-axis is the rank using whole-word representations.
+The y-axis is the rank using tokenized representations.
+Here, rank means the position of the expected word when all vocabulary words are sorted by similarity to the resulting embedding from vector arithmetic; lower is better.
+Points above the gray y=xy{=}x line mean tokenization hurts performance.
+Tokenization is an essential component of language modeling.
+To test how tokenization influences our findings, we use analogy completion tasks in English (e.g., man:king::woman:?) and compare two approaches: averaging subtoken embeddings after standard tokenization and summing embeddings from whole-word tokens.
+For each approach, we perform vector arithmetic on word representations (e.g., king - man + woman).
+We measure performance by ranking all vocabulary words by cosine similarity to the resulting representation, and observe how highly the expected word (e.g., queen) ranks, with a lower rank indicating better performance.
+Whole-word representations markedly outperform averaged subtokens across all models (Figure˜13), implying that linguistic regularities are primarily stored in whole-word embeddings rather than compositionally across subtokens.
+Despite tokenization effects, our classifier results show consistent patterns across models using different tokenizers (see Table˜5), indicating robust encoding of lexical and morphological information.
+Model HuggingFace ID
+BERT-Base bert-base-uncased
+BERT-Large bert-large-uncased
+DeBERTa-v3-Large microsoft/deberta-v3-large
+mT5-base google/mt5-base
+GPT-2-Small openai-community/gpt2
+GPT-2-Large openai-community/gpt2-large
+GPT-2-XL openai-community/gpt2-xl
+Pythia-6.9B EleutherAI/pythia-6.9b
+Pythia-6.9B-Tulu allenai/open-instruct-pythia-6.9b-tulu
+OLMo-2-1124-7B allenai/OLMo-2-1124-7B
+OLMo-2-1124-7B-Instruct allenai/OLMo-2-1124-7B-Instruct
+Gemma-2-2B google/gemma-2-2b
+Gemma-2-2B-Instruct google/gemma-2-2b-it
+Qwen2.5-1.5B Qwen/Qwen2.5-1.5B
+Qwen2.5-1.5B-Instruct Qwen/Qwen2.5-1.5B-Instruct
+Qwen2.5-7B Qwen/Qwen2.5-7B
+Qwen2.5-7B-Instruct Qwen/Qwen2.5-7B-Instruct
+Llama-3.1-8B meta-llama/Llama-3.1-8B
+Llama-3.1-8B-Instruct meta-llama/Llama-3.1-8B-Instruct
+Goldfish English goldfish-models/goldfish_eng_latn_1000mb
+Goldfish Chinese goldfish-models/goldfish_zho_hans_1000mb
+Goldfish German goldfish-models/goldfish_deu_latn_1000mb
+Goldfish French goldfish-models/goldfish_fra_latn_1000mb
+Goldfish Russian goldfish-models/goldfish_rus_cyrl_1000mb
+Goldfish Turkish goldfish-models/goldfish_tur_latn_1000mb
+Table 7: Canonical HuggingFace model IDs used to load models in our study.
+Appendix G Attention Head Analysis
+We conducted additional experiments analyzing attention head outputs alongside residual stream representations to understand how different components of transformer models contribute to linguistic encoding.
+G.1 Methodology
+We averaged activations across all attention heads at each layer for Qwen2.5-1.5B and Qwen2.5-1.5B-Instruct models using the English dataset.
+We then trained linear regression, MLP, and random forest classifiers on both attention head outputs and residual stream representations to compare their encoding patterns.
+G.2 Results
+Figure 14: Linguistic task accuracy for attention head outputs (solid lines) versus residual stream representations (dashed lines) across BERT and GPT-2 model families.
+Top row shows lemma accuracy, bottom row shows inflection accuracy.
+Columns represent different classifier types.
+Attention outputs show pronounced drops in lexical accuracy during middle layers, while residual streams maintain higher performance.
+Both streams maintain high inflectional accuracy, though residual streams consistently outperform attention outputs.
+Figure 15: Classifier selectivity for attention head outputs (solid lines) versus residual stream representations (dashed lines) across BERT and GPT-2 model families.
+Top row shows lemma selectivity, bottom row shows inflection selectivity.
+Attention outputs show near-zero lemma selectivity throughout all layers, while inflection selectivity reaches moderate levels (0.4-0.5) in both streams, with residual streams showing slightly higher values.
+Figure 16: Linguistic task accuracy for attention head outputs (solid lines) versus residual stream representations (dashed lines) across contemporary model families.
+Top row shows lemma accuracy, bottom row shows inflection accuracy.
+The pattern of attention outputs showing lower lexical accuracy in middle layers is consistent across all model families, while inflectional accuracy remains high in both streams.
+Figure 17: Classifier selectivity for attention head outputs (solid lines) versus residual stream representations (dashed lines) across contemporary model families.
+The selectivity patterns mirror those seen in BERT and GPT-2 families, with attention outputs maintaining near-zero lemma selectivity and moderate inflection selectivity across all models.
+The attention analysis reveals distinct encoding patterns between attention heads and residual streams.
+For lexical information, residual streams maintain relatively high accuracy (0.6-0.9) with gradual decline, while attention outputs show sharp drops to 0.2-0.4 in middle layers before modest recovery.
+For inflectional morphology, both components maintain high accuracy (0.7-1.0), though residual streams consistently outperform attention outputs, particularly in middle layers.
+Selectivity analysis confirms that attention heads encode inflectional information more than lexical information, with inflection selectivity reaching 0.4-0.5 while lemma selectivity remains near zero.
+This supports the hypothesis that attention mechanisms primarily handle contextual relationships rather than parametric lexical knowledge.
+Appendix H Steering Vector Analysis
+We conducted steering vector experiments to test whether inflectional representations can be functionally manipulated and to understand model sensitivity to activation interventions.
+H.1 Methodology
+For each inflectional category, we computed steering vectors as:
+$\mathbf{s}_{i}=\mathbf{\mu}_{i}-\lambda\cdot\frac{1}{|C|-1}\sum_{j\in C,j\neq i}\mathbf{\mu}_{j}$ (8)
+We tested multiple values of $\lambda$ (1, 5, 10, 20, 100) and measured the impact on MLP classifier performance when adding these steering vectors to existing activations for 1000 test words.
+We evaluated both mean probability change and prediction flip rate across all models.
+H.2 Results
+Figure 18: Mean probability change for inflection prediction when applying steering vectors across different $\lambda$ values.
+Five panels show results for $\lambda\in\{1,5,10,20,100\}$.
+All models start with high effectiveness ($\approx$0.9-1.0) at layer 0.
+Most models maintain stable performance, but Qwen2.5 variants show pronounced sensitivity dips around 10% layer depth before recovering.
+Higher $\lambda$ values increase steering effectiveness while preserving the overall pattern.
+Figure 19: Prediction flip rate when applying steering vectors across different $\lambda$ values.
+The flip rate patterns mirror the probability change results, with most models maintaining high rates (0.98-1.00) throughout all layers.
+Qwen2.5 variants show characteristic V-shaped dips to $\approx$0.60-0.70 around 10% layer depth.
+The consistency across $\lambda$ values suggests that steering effectiveness depends more on model architecture than intervention strength.
+Steering vector analysis demonstrates that inflectional representations can be effectively manipulated across most models and layers.
+The majority of models show robust steering effectiveness (>0.95 mean probability change and flip rate) throughout all layers, indicating that inflectional information is functionally accessible and modifiable.
+However, we observe model-specific sensitivity patterns.
+The Qwen2.5 variants show pronounced vulnerability in middle layers (around 10% depth), where steering effectiveness drops to 0.6-0.7 before recovering in later layers.
+This sensitivity pattern aligns with our intrinsic dimensionality findings, where these same models show extreme compression in middle layers (requiring only 1 component for 50-99% variance).
+The consistency of steering effectiveness across different $\lambda$ values suggests that the underlying representational structure, rather than intervention magnitude, determines steering success.
+This supports our conclusion that models with compressed middle-layer representations are more susceptible to activation interventions.
+Appendix I Random Forest Analysis
+For our in-depth analysis of lexical information and inflectional morphology, we evaluated Random Forest classifiers as an additional non-linear baseline to complement our linear regression and MLP analyses.
+I.1 Random Forest Classifier
+Random forests provide a complementary approach to linear models by handling complex feature interactions and decision boundaries.
+Unlike linear models that assume information is linearly separable, random forests combine an often large number of decision trees via model averaging to robustly identify patterns (Breiman, 2001).
+Since our classification tasks involve a large number of classes (thousands of unique lemmas and multiple inflection categories), we use a one-vs-all approach.
+For each class $j$, we train a binary random forest classifier that distinguishes class $j$ from all other classes.
+Each binary classifier outputs a probability score $s_{j}(x)\in[0,1]$ representing the confidence that input $x$ belongs to class $j$.
+The final prediction is assigned based on the maximum score across all classes:
+$\hat{y}=\operatorname{argmax}_{j\in\{1,2,...,c\}}s_{j}(x)$ (9)
+where $s_{j}(x)$ is the probability that the trained binary classifier assigns to class $j$ for the input xx [^5] We use scikit-learn’s implementation (Pedregosa et al., 2011) with the predict_proba() method to obtain probability scores.
+Figure 20: Random Forest classifier performance across model layers for English.
+The left column shows accuracy for lemma (top) and inflection (bottom) prediction.
+The right column shows classifier selectivity (difference between linguistic and control task accuracy) for the same tasks.
+Each line represents a different model.
+Figure 21: Random Forest classifier performance across languages and model layers.
+The left column shows accuracy for inflection prediction (lemma prediction omitted due to computational constraints with thousands of classes).
+The right column shows classifier selectivity.
+I.2 English Results
+Lexical Identity
+Random Forest classifiers for lemma prediction were not comprehensively evaluated due to computational constraints arising from thousands of unique lemma classes.
+The one-vs-all training approach required for multi-class classification with Random Forests becomes prohibitively expensive with such a large number of classes.
+Inflectional Morphology
+Random Forest classifiers for inflectional morphology show markedly different patterns compared to linear regression and MLP classifiers (Figure˜20, bottom left).
+While linear and MLP classifiers maintain consistently high accuracy (0.9-1.0) across all layers, Random Forests exhibit lower overall accuracy (0.65-0.90) and a declining trend across layers.
+This contrasts sharply with the flat, high-performance curves observed for linear methods.
+Selectivity Analysis
+Random Forest selectivity for inflectional morphology (Figure˜20, bottom right) shows moderate values (0.2-0.4) but with more variability across models compared to linear methods.
+This suggests that while Random Forests can generalize beyond memorization, they are less consistent in their ability to extract the underlying morphological patterns compared to simpler linear approaches.
+I.3 Cross-Linguistic Patterns
+The multilingual Random Forest analysis (Figure˜21) reveals consistent cross-linguistic patterns while highlighting language-specific challenges.
+Cross-Linguistic Performance
+Random Forest classifiers for inflectional morphology show the same language ordering observed with linear and MLP classifiers: Russian and German achieve the highest accuracy (0.80-0.90), English and French show intermediate performance (0.70-0.85), Chinese maintains moderate accuracy (0.65-0.80), and Turkish consistently underperforms (0.55-0.75).
+This consistency across classifier types confirms that the observed patterns reflect genuine differences in how linguistic information is encoded rather than classifier-specific artifacts.
+Selectivity Patterns
+Random Forest selectivity patterns are similar to those of linear classifiers but with greater variance.
+Russian and German maintain the highest selectivity (0.3-0.5), while Turkish shows the lowest (0.1-0.3).
+This consistency suggests that the underlying representational structure, rather than classifier choice, drives the observed cross-linguistic differences.
+Appendix J Classifier Error Analysis
+We conducted a detailed error analysis of our classifiers to better understand their performance across different morphological features and languages.
+See Table˜8 through Table˜26 for the full results.
 Model 3rd person Base Comparative Past Plural Positive Singular Superlative
@@
 Qwen2.5-1.5B-Instruct 0.774 0.954 0.647 0.945 0.972 0.921 0.965 0.630
-Table 8: Breakdown of inflection classification accuracy by morphological feature for each model, using linear regression classifiers.
+Table 8: Breakdown of inflection classification accuracy by morphological feature for each model using linear regression classifiers (English).
 Inflections are grouped by their morphological features (e.g., Past, Plural, Comparative).
@@
 All accuracy values are on a 0–1 scale.
+Comparative and superlative forms consistently show the lowest accuracy across all models, reflecting the challenges of these less frequent morphological categories.
 Model 3rd person Base Comparative Past Plural Positive Singular Superlative
@@
 Qwen2.5-1.5B-Instruct 0.808 0.953 0.697 0.947 0.974 0.930 0.965 0.682
-Table 9: Breakdown of inflection classification accuracy by morphological feature for each model, using MLP classifiers.
+Table 9: Breakdown of inflection classification accuracy by morphological feature for each model using MLP classifiers (English).
 Inflections are grouped by their morphological features (e.g., Past, Plural, Comparative).
@@
 All accuracy values are on a 0–1 scale.
+MLP classifiers provide modest improvements over linear regression, particularly for comparative and superlative forms, though the relative ordering across morphological features remains consistent.
 Model Noun Verb Adjective Adverb Pronoun Preposition Conjunction Interjection Other
@@
 Qwen2.5-1.5B-Instruct 0.642 0.800 0.632 0.831 0.544 0.082 0.245 0.068 0.877
-Llama-3-8B 0.776 0.882 0.771 0.887 0.831 0.286 0.396 0.321 0.911
-Llama-3-8B-Instruct 0.796 0.892 0.788 0.896 0.908 0.300 0.443 0.357 0.917
-Table 10: Breakdown of lexeme classification accuracy by Part of Speech (POS) for each model, using linear regression classifiers.
-Lexemes are grouped by their POS tags (e.g., Noun, Verb, Adjective).
+Llama-3.1-8B 0.776 0.882 0.771 0.887 0.831 0.286 0.396 0.321 0.911
+Llama-3.1-8B-Instruct 0.796 0.892 0.788 0.896 0.908 0.300 0.443 0.357 0.917
+Table 10: Breakdown of lemma classification accuracy by Part of Speech (POS) for each model using linear regression classifiers (English).
+Lemmas are grouped by their POS tags (e.g., Noun, Verb, Adjective).
 For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer.
 All accuracy values are on a 0–1 scale.
+Performance varies significantly with frequency: frequent categories like nouns and verbs achieve higher accuracy, while infrequent categories like pronouns and prepositions show lower performance due to limited training examples.
 Model Noun Verb Adjective Adverb Pronoun Preposition Conjunction Interjection Other
@@
 Qwen2.5-1.5B-Instruct 0.749 0.840 0.738 0.818 0.211 0.098 0.564 0.123 0.860
-Llama-3-8B 0.798 0.879 0.807 0.886 0.800 0.214 0.679 0.393 0.882
-Llama-3-8B-Instruct 0.824 0.893 0.826 0.895 0.831 0.257 0.689 0.429 0.887
-Table 11: Breakdown of lexeme classification accuracy by Part of Speech (POS) for each model, using Multi-Layer Perceptron (MLP) classifiers.
-Lexemes are grouped by their POS tags (e.g., Noun, Verb, Adjective).
+Llama-3.1-8B 0.798 0.879 0.807 0.886 0.800 0.214 0.679 0.393 0.882
+Llama-3.1-8B-Instruct 0.824 0.893 0.826 0.895 0.831 0.257 0.689 0.429 0.887
+Table 11: Breakdown of lemma classification accuracy by Part of Speech (POS) for each model using Multi-Layer Perceptron (MLP) classifiers (English).
+Lemmas are grouped by their POS tags (e.g., Noun, Verb, Adjective).
 For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer.
+All accuracy values are on a 0–1 scale.
+MLP classifiers provide consistent improvements over linear regression across all POS categories, though the frequency-dependent performance patterns persist.
+Model Linear Regression MLP
+Positive Base Plural Singular Positive Base Plural Singular
+(n=300) (n=2,074) (n=3) (n=3,947) (n=300) (n=2,074) (n=3) (n=3,947)
+mT5-Base 0.739 0.913 0.436 0.962 0.783 0.919 0.231 0.961
+Qwen2.5-1.5B 0.785 0.929 0.034 0.969 0.801 0.924 0.092 0.967
+Qwen2.5-1.5B-Instruct 0.779 0.925 0.034 0.964 0.803 0.923 0.057 0.967
+Qwen2.5-7B 0.824 0.937 0.310 0.970 0.828 0.929 0.310 0.969
+Qwen2.5-7B-Instruct 0.819 0.936 0.299 0.970 0.823 0.928 0.276 0.969
+Goldfish Chinese 0.793 0.912 0.000 0.958 0.816 0.915 0.000 0.957
+Table 12: Breakdown of inflection classification accuracy for each model by inflection type using Linear Regression and Multi-Layer Perceptron (MLP) classifiers (Chinese).
+Accuracies are calculated over all examples for a given group across all layers.
+Counts (n) are derived from a single representative layer for each group.
+All accuracy values are on a 0–1 scale.
+Model Noun Verb Adjective Adverb Preposition Other
+(n=1,179) (n=564) (n=108) (n=22) (n=20) (n=50)
+mT5-Base 0.838 0.828 0.786 0.762 0.920 0.726
+Qwen2.5-1.5B 0.810 0.797 0.746 0.715 0.872 0.699
+Qwen2.5-1.5B-Instruct 0.813 0.799 0.748 0.713 0.873 0.700
+Qwen2.5-7B 0.887 0.882 0.846 0.847 0.915 0.817
+Qwen2.5-7B-Instruct 0.886 0.877 0.843 0.835 0.913 0.811
+Goldfish Chinese 0.883 0.878 0.845 0.875 0.954 0.858
+Table 13: Breakdown of lemma classification accuracy by Part of Speech (POS) for each model, using Linear Regression classifiers (Chinese).
+Lemmas are grouped by their POS tags (e.g., , Noun, Verb, Adjective).
+Accuracies are calculated over all examples for a given group across all layers.
+Counts (n) are derived from a single representative layer for each group.
+All accuracy values are on a 0–1 scale.
+Model Noun Verb Adjective Adverb Preposition Other
+(n=1,179) (n=564) (n=108) (n=22) (n=20) (n=50)
+mT5-Base 0.698 0.712 0.564 0.571 0.884 0.569
+Qwen2.5-1.5B 0.748 0.761 0.658 0.668 0.826 0.669
+Qwen2.5-1.5B-Instruct 0.735 0.745 0.643 0.643 0.814 0.655
+Qwen2.5-7B 0.815 0.826 0.749 0.745 0.848 0.750
+Qwen2.5-7B-Instruct 0.815 0.822 0.747 0.734 0.845 0.744
+Goldfish Chinese 0.766 0.771 0.647 0.621 0.912 0.682
+Table 14: Breakdown of lemma classification accuracy by Part of Speech (POS) for each model, using Multi-Layer Perceptron (MLP) classifiers (Chinese).
+Lemmas are grouped by their POS tags (e.g., , Noun, Verb, Adjective).
+Accuracies are calculated over all examples for a given group across all layers.
+Counts (n) are derived from a single representative layer for each group.
+All accuracy values are on a 0–1 scale.
+Model Base 3rd person Positive Past Plural Superlative Singular Comparative
+(n=417) (n=517) (n=1,720) (n=839) (n=1,076) (n=52) (n=3,197) (n=141)
+mT5-Base 0.908 0.941 0.940 0.960 0.882 0.572 0.962 0.636
+Qwen2.5-1.5B 0.849 0.889 0.922 0.914 0.888 0.657 0.953 0.796
+Qwen2.5-1.5B-Instruct 0.844 0.887 0.922 0.910 0.889 0.659 0.952 0.795
+Qwen2.5-7B 0.892 0.922 0.939 0.947 0.909 0.826 0.962 0.878
+Qwen2.5-7B-Instruct 0.915 0.934 0.945 0.962 0.924 0.866 0.968 0.909
+Goldfish German 0.938 0.941 0.955 0.979 0.916 0.542 0.968 0.708
+Table 15: Breakdown of inflection classification accuracy for each model by inflection type using Linear Regression classifiers (German).
+Accuracies are calculated over all examples for a given group across all layers.
+Counts (n) are derived from a single representative layer for each group.
+All accuracy values are on a 0–1 scale.
+Model Base 3rd person Positive Past Plural Superlative Singular Comparative
+(n=417) (n=517) (n=1,720) (n=839) (n=1,076) (n=52) (n=3,197) (n=141)
+mT5-Base 0.921 0.945 0.948 0.959 0.884 0.723 0.967 0.770
+Qwen2.5-1.5B 0.890 0.915 0.930 0.940 0.897 0.831 0.958 0.892
+Qwen2.5-1.5B-Instruct 0.888 0.914 0.930 0.938 0.898 0.825 0.957 0.897
+Qwen2.5-7B 0.912 0.932 0.944 0.956 0.913 0.868 0.964 0.924
+Qwen2.5-7B-Instruct 0.925 0.941 0.950 0.966 0.928 0.901 0.970 0.936
+Goldfish German 0.947 0.957 0.964 0.978 0.923 0.817 0.970 0.896
+Table 16: Breakdown of inflection classification accuracy for each model by inflection type using Multi-Layer Perceptron (MLP) classifiers (German).
+Accuracies are calculated over all examples for a given group across all layers.
+Counts (n) are derived from a single representative layer for each group.
+All accuracy values are on a 0–1 scale.
+Model Linear Regression MLP
+Noun Verb Adjective Other Noun Verb Adjective Other
+(n=1,262) (n=395) (n=406) (n=12) (n=1,262) (n=395) (n=406) (n=12)
+mT5-Base 0.685 0.662 0.568 0.750 0.611 0.602 0.486 0.723
+Qwen2.5-1.5B 0.743 0.725 0.715 0.775 0.721 0.700 0.687 0.711
+Qwen2.5-1.5B-Instruct 0.740 0.722 0.715 0.766 0.722 0.698 0.687 0.704
+Qwen2.5-7B 0.821 0.809 0.808 0.829 0.795 0.786 0.783 0.814
+Qwen2.5-7B-Instruct 0.815 0.803 0.803 0.821 0.795 0.785 0.782 0.813
+Goldfish German 0.720 0.747 0.701 0.769 0.758 0.772 0.742 0.769
+Table 17: Breakdown of lemma classification accuracy by Part of Speech (POS) for each model, using Linear Regression and Multi-Layer Perceptron (MLP) classifiers (German).
+Lemmas are grouped by their POS tags (e.g., , Noun, Verb, Adjective).
+Accuracies are calculated over all examples for a given group across all layers.
+Counts (n) are derived from a single representative layer for each group.
+All accuracy values are on a 0–1 scale.
+Model Base 3rd person Positive Past Plural Singular
+(n=688) (n=776) (n=1,833) (n=857) (n=1,457) (n=5,169)
+mT5-Base 0.934 0.912 0.879 0.908 0.954 0.970
+Qwen2.5-1.5B 0.933 0.858 0.896 0.903 0.958 0.967
+Qwen2.5-1.5B-Instruct 0.930 0.852 0.893 0.898 0.958 0.966
+Qwen2.5-7B 0.955 0.918 0.918 0.931 0.965 0.975
+Qwen2.5-7B-Instruct 0.951 0.913 0.915 0.928 0.964 0.974
+Goldfish French 0.942 0.955 0.937 0.930 0.968 0.976
+Table 18: Breakdown of inflection classification accuracy for each model by inflection type using Linear Regression classifiers (French).
+Accuracies are calculated over all examples for a given group across all layers.
+Counts (n) are derived from a single representative layer for each group.
+All accuracy values are on a 0–1 scale.
+Model Base 3rd person Positive Past Plural Singular
+(n=688) (n=776) (n=1,833) (n=857) (n=1,457) (n=5,169)
+mT5-Base 0.957 0.937 0.911 0.935 0.957 0.977
+Qwen2.5-1.5B 0.954 0.905 0.914 0.925 0.965 0.968
+Qwen2.5-1.5B-Instruct 0.954 0.902 0.911 0.924 0.965 0.968
+Qwen2.5-7B 0.966 0.936 0.930 0.937 0.970 0.976
+Qwen2.5-7B-Instruct 0.962 0.931 0.926 0.934 0.970 0.975
+Goldfish French 0.974 0.967 0.945 0.942 0.973 0.979
+Table 19: Breakdown of inflection classification accuracy for each model by inflection type using Multi-Layer Perceptron (MLP) classifiers (French).
+Accuracies are calculated over all examples for a given group across all layers.
+Counts (n) are derived from a single representative layer for each group.
+All accuracy values are on a 0–1 scale.
+Model Linear Regression MLP
+Noun Verb Adjective Other Noun Verb Adjective Other
+(n=1,496) (n=406) (n=358) (n=15) (n=1,496) (n=406) (n=358) (n=15)
+mT5-Base 0.708 0.577 0.605 0.799 0.755 0.560 0.636 0.820
+Qwen2.5-1.5B 0.754 0.725 0.673 0.824 0.807 0.765 0.751 0.853
+Qwen2.5-1.5B-Instruct 0.750 0.718 0.671 0.820 0.824 0.776 0.768 0.869
+Qwen2.5-7B 0.840 0.814 0.764 0.869 0.856 0.825 0.794 0.884
+Qwen2.5-7B-Instruct 0.833 0.805 0.758 0.860 0.851 0.818 0.792 0.883
+Goldfish French 0.749 0.758 0.661 0.811 0.894 0.869 0.813 0.888
+Table 20: Breakdown of lemma classification accuracy by Part of Speech (POS) for each model, using Linear Regression and Multi-Layer Perceptron (MLP) classifiers (French).
+Lemmas are grouped by their POS tags (e.g., , Noun, Verb, Adjective).
+Accuracies are calculated over all examples for a given group across all layers.
+Counts (n) are derived from a single representative layer for each group.
+All accuracy values are on a 0–1 scale.
+Model Base 3rd person Positive Past Plural Superlative Singular Comparative
+(n=690) (n=456) (n=1,192) (n=455) (n=1,333) (n=3) (n=3,316) (n=23)
+mT5-Base 0.930 0.978 0.975 0.957 0.877 0.000 0.977 0.799
+Qwen2.5-1.5B 0.925 0.946 0.974 0.938 0.923 0.015 0.966 0.835
+Qwen2.5-1.5B-Instruct 0.924 0.943 0.974 0.934 0.921 0.015 0.966 0.817
+Qwen2.5-7B 0.949 0.966 0.979 0.958 0.948 0.094 0.977 0.872
+Qwen2.5-7B-Instruct 0.951 0.974 0.980 0.970 0.948 0.080 0.980 0.918
+Goldfish Russian 0.940 0.950 0.976 0.931 0.921 0.000 0.976 0.867
+Table 21: Breakdown of inflection classification accuracy for each model by inflection type using Linear Regression classifiers (Russian).
+Accuracies are calculated over all examples for a given group across all layers.
+Counts (n) are derived from a single representative layer for each group.
+All accuracy values are on a 0–1 scale.
+Model Base 3rd person Positive Past Plural Superlative Singular Comparative
+(n=690) (n=456) (n=1,192) (n=455) (n=1,333) (n=3) (n=3,316) (n=23)
+mT5-Base 0.959 0.978 0.969 0.966 0.904 0.000 0.978 0.849
+Qwen2.5-1.5B 0.952 0.955 0.972 0.948 0.933 0.089 0.970 0.899
+Qwen2.5-1.5B-Instruct 0.950 0.954 0.973 0.947 0.933 0.089 0.969 0.911
+Qwen2.5-7B 0.963 0.964 0.978 0.960 0.951 0.246 0.979 0.910
+Qwen2.5-7B-Instruct 0.961 0.970 0.978 0.966 0.949 0.126 0.980 0.924
+Goldfish Russian 0.965 0.972 0.978 0.948 0.943 0.000 0.977 0.934
+Table 22: Breakdown of inflection classification accuracy for each model by inflection type using Multi-Layer Perceptron (MLP) classifiers (Russian).
+Accuracies are calculated over all examples for a given group across all layers.
+Counts (n) are derived from a single representative layer for each group.
+All accuracy values are on a 0–1 scale.
+Model Linear Regression MLP
+Noun Verb Adjective Other Noun Verb Adjective Other
+(n=982) (n=333) (n=275) (n=4) (n=982) (n=333) (n=275) (n=4)
+mT5-Base 0.660 0.614 0.542 0.648 0.492 0.484 0.387 0.426
+Qwen2.5-1.5B 0.777 0.712 0.759 0.720 0.712 0.696 0.716 0.647
+Qwen2.5-1.5B-Instruct 0.772 0.704 0.756 0.720 0.710 0.689 0.717 0.643
+Qwen2.5-7B 0.854 0.790 0.843 0.812 0.798 0.794 0.813 0.749
+Qwen2.5-7B-Instruct 0.845 0.778 0.835 0.807 0.794 0.785 0.809 0.744
+Goldfish Russian 0.795 0.723 0.764 0.676 0.810 0.776 0.759 0.657
+Table 23: Breakdown of lemma classification accuracy by Part of Speech (POS) for each model, using Linear Regression and Multi-Layer Perceptron (MLP) classifiers (Russian).
+Lemmas are grouped by their POS tags (e.g., , Noun, Verb, Adjective).
+Accuracies are calculated over all examples for a given group across all layers.
+Counts (n) are derived from a single representative layer for each group.
+All accuracy values are on a 0–1 scale.
+Model Base 3rd person Positive Past Plural Singular
+(n=154) (n=51) (n=401) (n=168) (n=33) (n=632)
+mT5-Base 0.860 0.911 0.928 0.966 0.837 0.952
+Qwen2.5-1.5B 0.808 0.802 0.721 0.928 0.861 0.892
+Qwen2.5-1.5B-Instruct 0.809 0.817 0.720 0.941 0.878 0.899
+Qwen2.5-7B 0.865 0.879 0.810 0.966 0.903 0.909
+Qwen2.5-7B-Instruct 0.850 0.874 0.796 0.960 0.886 0.900
+Goldfish Turkish 0.847 0.915 0.880 0.964 0.872 0.963
+Table 24: Breakdown of inflection classification accuracy for each model by inflection type using Linear Regression classifiers (Turkish).
+Accuracies are calculated over all examples for a given group across all layers.
+Counts (n) are derived from a single representative layer for each group.
+All accuracy values are on a 0–1 scale.
+Model Base 3rd person Positive Past Plural Singular
+(n=154) (n=51) (n=401) (n=168) (n=33) (n=632)
+mT5-Base 0.755 0.760 0.848 0.922 0.515 0.949
+Qwen2.5-1.5B 0.770 0.767 0.667 0.919 0.765 0.914
+Qwen2.5-1.5B-Instruct 0.762 0.757 0.662 0.917 0.766 0.913
+Qwen2.5-7B 0.853 0.845 0.791 0.956 0.875 0.937
+Qwen2.5-7B-Instruct 0.845 0.844 0.786 0.956 0.875 0.932
+Goldfish Turkish 0.832 0.879 0.870 0.957 0.834 0.957
+Table 25: Breakdown of inflection classification accuracy for each model by inflection type using Multi-Layer Perceptron (MLP) classifiers (Turkish).
+Accuracies are calculated over all examples for a given group across all layers.
+Counts (n) are derived from a single representative layer for each group.
+All accuracy values are on a 0–1 scale.
+Model Linear Regression MLP
+Noun Verb Adjective Other Noun Verb Adjective Other
+(n=221) (n=53) (n=104) (n=13) (n=221) (n=53) (n=104) (n=13)
+mT5-Base 0.866 0.823 0.921 0.955 0.215 0.421 0.374 0.637
+Qwen2.5-1.5B 0.834 0.805 0.866 0.877 0.307 0.439 0.449 0.693
+Qwen2.5-1.5B-Instruct 0.816 0.791 0.860 0.874 0.305 0.439 0.448 0.691
+Qwen2.5-7B 0.871 0.850 0.900 0.904 0.595 0.625 0.695 0.809
+Qwen2.5-7B-Instruct 0.850 0.823 0.883 0.885 0.579 0.613 0.678 0.800
+Goldfish Turkish 0.929 0.904 0.940 0.969 0.386 0.550 0.477 0.808
+Table 26: Breakdown of lemma classification accuracy by Part of Speech (POS) for each model, using Linear Regression and Multi-Layer Perceptron (MLP) classifiers (Turkish).
+Lemmas are grouped by their POS tags (e.g., , Noun, Verb, Adjective).
+Accuracies are calculated over all examples for a given group across all layers.
+Counts (n) are derived from a single representative layer for each group.
 All accuracy values are on a 0–1 scale.
@@
 (b) Linear separability gap for lemma prediction
-Figure 7: Performance advantage of MLP classifiers over linear classifiers (in percentage points) across model layers.
+Figure 22: Performance advantage of MLP classifiers over linear classifiers (in percentage points) across model layers for English.
 The linear separability gap measures how much a non-linear transformation improves classifier performance compared to a simple linear mapping.
-For inflection prediction, the gap is consistently minimal (mostly within ±$\pm 0.02$ percentage points) and sometimes negative, indicating that inflectional features are primarily encoded in a linear fashion throughout the network.
+For inflection prediction, the gap is consistently minimal (mostly within $\pm 0.02$ percentage points) and sometimes negative, indicating that inflectional features are primarily encoded in a linear fashion throughout the network.
 By contrast, the linear separability gap for lemma prediction is relatively large (0.1–0.3 percentage points) and positive across most models
-Figure 8: Intrinsic dimensionality curves for all 10 models.
-Each subplot shows the relationship between the percentage of maximum PCA components (x-axis) and the percentage of explained variance (y-axis) across different layers.
-The color gradient from purple (early layers, 0%) to yellow (late layers, 100%) indicates the relative layer depth within each model.
-Models like BERT, Gemma, and Llama show similar compression patterns, while GPT-2 variants, Qwen and Pythia exhibit opposite trends in their middle layers.
-Figure 9: Lexeme (top row) and inflection (bottom row) prediction accuracy across normalized layer number for OLMo-2-7B checkpoints at various pretraining steps (5k-734k steps).
+Cross-Linguistic Linear Separability Gap
+(a) Chinese - Inflection
+(b) Chinese - Lemma
+(c) German - Inflection
+(d) German - Lemma
+(e) French - Inflection
+(f) French - Lemma
+(g) Russian - Inflection
+(h) Russian - Lemma
+(i) Turkish - Inflection
+(j) Turkish - Lemma
+Figure 23: Cross-linguistic linear separability gap showing performance advantage of MLP classifiers over linear classifiers across model layers for five additional languages.
+For inflectional features, mT5 and Goldfish models show slight positive gaps (indicating modest benefits from non-linear classification), while Qwen2.5 variants show slight negative gaps (indicating linear classifiers are sufficient or superior).
+For lexical features, all models show negative gaps that are most pronounced in early layers, suggesting that linear regression with regularization consistently outperforms MLPs for lexical classification across all model families and languages.
+Figure 24: Lemma (top row) and inflection (bottom row) prediction accuracy across normalized layer number for OLMo-2-7B checkpoints at various pretraining steps (5k-734k steps) for English.
 The full model is 928k steps.
 Checkpoints are on a color gradient from brightest (earliest) to darkest (latest).
-Early checkpoints exhibit higher lexeme accuracy than later ones, while inflectional accuracy remains flat across layers and checkpoints.
-Figure 10: Lexeme (top row) and inflection (bottom row) prediction accuracy across normalized layer number for Pythia-6.9B checkpoints at various pretraining steps (1-111k steps).
+Early checkpoints exhibit higher lemma accuracy than later ones, while inflectional accuracy remains flat across layers and checkpoints.
+Figure 25: Lemma (top row) and inflection (bottom row) prediction accuracy across normalized layer number for Pythia-6.9B checkpoints at various pretraining steps (1-111k steps) for English.
 The full model is 143k steps.
 Checkpoints are on a color gradient from brightest (earliest) to darkest (latest).
-Lexeme accuracy declines both with deeper layers and with more training, whereas inflectional accuracy stays uniformly high across all layers and checkpoints.
+Lemma accuracy declines both with deeper layers and with more training, whereas inflectional accuracy stays uniformly high across all layers and checkpoints.